### v4

# Dynamical systems defined by polynomials with algebraic properties

Abstract
Let $St(\mathbb{T})$ be the set of streams, namely the two side sequences whose terms take value from the torus $\mathbb{T}$. We are mainly interested in the algebraic and dynamical properties of the space of stream zeros $\Omega_{P}$, which is the kernel via convolution product by a primitive polynomial $P(z)\in\mathbb{Z}[z]$ acting on $St(\mathbb{T})$. Together with the shift $\sigma$, it defines a topological dynamics $(\Omega_{P},\sigma)$ whose symbolic representation is given by the integer part of the action. It is also proved that two primitive polynomials $P(z),Q(z)$ are coprime if and only if $\dim(\Omega_{P}\cap\Omega_{Q})=0$. Furthermore, the group of (strong) automorphisms of $\Omega_{P}$ commuting with the shift $\sigma$ is also investigated. Specially in the case of $\deg P=2$, it is determined by the discriminant of equation $P(z)=0$. Finally, we study the streams over the Galois field $GF(q^{k})=GF(q)(\theta)$, where $\theta$ is a root of an irreducible polynomial $P(z)\in GF(q)[z]$ of degree $k$. The zeros $\Omega_{P}^{q}$ of $ST(GF(q^{k}))$ by the action of $P(z)$ in the above sense is discussed. In this case, the group of automorphisms of $\Omega_{P}^{q}$ commuting with the shift is proved to be a cyclic group of order $q^{k}-1$.
1 Introduction
Let $A=(a_{n};n\in{Z})$ be a two-side sequence of elements in an additive group $G$, which is called a two side stream over $G$. The set of two side streams over $G$ is denoted by $St(G)$. This $A$ is written as $(\cdots,a_{-2},a_{-1};a_{0},a_{1},\cdots)$ and is also viewed as a two-sided formal series on a variable,say $z$, namely,
$A=\cdots+a_{-2}z^{-2}+a_{-1}z^{-1}+a_{0}+a_{1}z+a_{2}z^{2}+\cdots.$
Note that $St(G)$ is an additive group by component-wise addition. Moreover, if $G$ is a topological group, then $St(G)$ is considered as a topological group with the product topology coming from $G$. Let 00 be the additive identity of $G$, then the additive identity of $St(G)$ is $(0^{\infty};0^{\infty})$, where $0^{\infty}$ before “;” implies $\cdots,0,0$ while that after “;” implies $0,0,\cdots$.
Let $G$ be a topological commutative ring having the multiplication identity 1. We define the convolution product of $(a_{n};n\in\mathbb{Z})$ and $(b_{n};n\in\mathbb{Z})\in St(G),$ if the following infinite series exists:
$(a_{n};n\in\mathbb{Z})\times(b_{n};n\in\mathbb{Z})=\left(\sum_{i=-\infty}^{\infty}a_{i}b_{n-i};~{}n\in\mathbb{Z}\right).$
The identity of $St(G)$ with respect to the convolution product is $I=(0^{\infty};1,0^{\infty})$. We call $(a_{n};n\in\mathbb{Z})\in St(G)$ absolutely summable if the series $\sum_{n\in\mathbb{Z}}a_{n}$ converges to the same value that is independent of the order of the sum.
One-sided sequences of $G$ have been studied for a long time in the field of computer science as stream calculus (see [2], for example), where the convolution product always exists since the above sum becomes a finite sum. In this paper, we study two side sequences and call them streams.
For $(a_{n};n\in\mathbb{Z})\in St(G)$, even if $(b_{n};~{}n\in\mathbb{Z})\in St(G)$ such that
$(a_{n};n\in\mathbb{Z})\times(b_{n};~{}n\in\mathbb{Z})=I$ (1)
exists, it may not be unique. For example,
$(0^{\infty};1,-1,0^{\infty})\times(0^{\infty};1^{\infty})=(0^{\infty};1,-1,0^{\infty})\times((-1)^{\infty};0^{\infty})=I.$
An element $(a_{n};n\in\mathbb{Z})$ of $St(\mathbb{C})$ such that $a_{n}=0$ except for finitely many $n$ can be identified with a polynomial $P(z)=\sum_{n\in S}a_{n}z^{n}$ admitting negative powers, where $S\subset\mathbb{Z}$ is the finite set of $n$ such that $a_{n}\neq 0$. We always assume that $P(z)$ is nontrivial, that is, $S\neq\emptyset$. We call $P(z)$ hyperbolic if $P(z)=0$ does not hold for any $z\in\mathbb{C}$ with $|z|=1$.
We are especially interested in polynomials with integer coefficients whose greatest common divisor (GCD) is 1. We admit polynomials with negative power, but in most cases, it does not lose generality to restrict them only with nonnegative powers, so we will do sometimes.
Let $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ be the torus, $P(z)$ acts on $St(\mathbb{T})$ by the convolution product in the following way
$(P(z),(x_{n};n\in\mathbb{Z}))\mapsto(a_{n};n\in S)\times(x_{n};n\in\mathbb{Z})\in St(\mathbb{T}).$
We study the kernel of this action, denoted by $\Omega_{P}$ and is called the stream zeros of $P(z)$, that is
$\Omega_{P}:=\left\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T}):P(z)\times(x_{n};n\in\mathbb{Z})=(0^{\infty};0^{\infty})\right\}.$
We show that $\Omega_{P}$ has a lot of similarities with the algebraic roots of $P(z)=0$. This might provide us with a way of understanding the roots of algebraic equations.
Let $GF(q)$ be the Galois field of size $q$, where $q$ is a power of a prime. Let $P(z)$ be a polynomial in $GF(q)[z]$. In the last section, we discuss the zeros of $P(z)$-action to $St(GF(q))$, which is denoted by $\Omega_{P}^{q}$.
2 Convolution product and its inverse
In this section, we present some basic properties of the convolution product of $St(\mathbb{C})$. As one can see from the previous discussion, the convolution inverse may not be unique, we thus wish to find some conditions that can guarantee the uniqueness of the convolution inverse. The following theorem is known ([4]), here we give another elementary proof.
Theorem 1.
(i) Among $A,B,C\in St(\mathbb{C})$, two of them are absolutely summable and the remaining one is bounded, then the convolution products below exist and it holds that
$(A\times B)\times C=A\times(B\times C).$
(ii) For any absolutely summable $(a_{n};n\in\mathbb{Z})\in St(\mathbb{C})$, an absolutely summable $(b_{n};n\in\mathbb{Z})$ satisfying $(1)$ is unique if it exists. We call $(b_{n};n\in\mathbb{Z})$ as this convolution inverse of $(a_{n};n\in\mathbb{Z})$ and is denoted by $(a_{n};n\in\mathbb{Z})^{-1}$.
(iii) If both of $(a_{n};n\in\mathbb{Z}),~{}(b_{n};n\in\mathbb{Z})\in St(\mathbb{C})$ are absolutely summable, then so is their convolution product. Furthermore, if both of $(a_{n};n\in\mathbb{Z})^{-1},(b_{n};n\in\mathbb{Z})^{-1}$ exist, then
$\left((a_{n};n\in\mathbb{Z})\times(b_{n};n\in\mathbb{Z})\right)^{-1}$
exists as well and
$\left((a_{n};n\in\mathbb{Z})\times(b_{n};n\in\mathbb{Z})\right)^{-1}=(a_{n};n\in\mathbb{Z})^{-1}\times(b_{n};n\in\mathbb{Z})^{-1}$
holds. Moreover, we have
$\left((a_{n};n\in\mathbb{Z})^{-1}\right)^{-1}=(a_{n};n\in\mathbb{Z}).$
(iv) Any hyperbolic polynomial $P=(a_{n};n\in\mathbb{Z})\not\equiv 0$ ($P=(a_{n};n\in\mathbb{Z})\neq(0^{\infty};0^{\infty})$ which is better? has an (comment: ”a unique” seems to be better)absolutely summable inverse.
Proof.
(i) Let
$A=(a_{n};n\in\mathbb{Z}),~{}B=(b_{n};n\in\mathbb{Z}),~{}C=(c_{n};n\in\mathbb{Z}).$
Then for any $n\in\mathbb{Z}$, we prove that
$\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}|a_{i}b_{j}c_{k}|<\infty.$ (2)
By symmetry, we may assume that $A,B$ are absolutely summable and $C$ is bounded. Then, there exists a constant $L>0$ such that
$\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}|a_{i}b_{j}c_{k}|<L\sum_{i,j\in\mathbb{Z}}|a_{i}b_{j}|=L\sum_{i\in\mathbb{Z}}|a_{i}|\sum_{j\in\mathbb{Z}}|b_{j}|<\infty.$
Thus, $\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}a_{i}b_{j}c_{k}$ converges absolutely. Hence, we have
$\sum_{k\in\mathbb{Z}}\left(\sum_{i,j\in\mathbb{Z}\atop{i+j=n-k}}a_{i}b_{j}\right)c_{k}=\sum_{i\in\mathbb{Z}}a_{i}\left(\sum_{j,k\in\mathbb{Z}\atop{j+k=n-i}}b_{j}c_{k}\right)=\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}a_{i}b_{j}c_{k}$
and $(A\times B)\times C=A\times(B\times C)$.
(ii) Suppose that any of $A,B,C\in St(\mathbb{C})$ be absolutely summable and both of $B,C$ are convolution inverses of $A$. Then by (i), we have $B=B\times(A\times C)=(B\times A)\times C=C$. Thus, the absolutely summable convolution inverse of $A$ is unique.
(iii) is clear.
(iv) Let $P(z)=\sum_{n\in S}a_{n}z^{n}$ with a finite nonempty set $S=\{n\in\mathbb{Z};a_{n}\neq 0\}$. Consider the factorization of $P(z)$ and applying (ii), it is sufficient to prove that $P(z)$ being of the form, i.e., constant $C\neq 0$, $z,z^{-1}$ and $z-\omega$ with $0<|\omega|\neq 1$, has the absolutely summable inverse.
Case 1: $P(z)=C$. Since $C=(0^{\infty};C,0^{\infty})$, $C^{-1}=(0^{\infty};c^{-1},0^{\infty})$ becomes the absolutely summable inverse.
Case 2: $P(z)=z~{}\mbox{or}~{}z^{-1}$. One easily verifies that $z$ and $z^{-1}$ are absolutely summable and inverses of each other. That is, $(0^{\infty}1;0^{\infty})=(0^{\infty};0,1,0^{\infty})^{-1}$.
Case 3: $P(z)=z-w$ with $|\omega|>1.$ Since
$(z-\omega)^{-1}=(-1)\omega^{-1}(1-\omega^{-1}z)^{-1}=-\omega^{-1}-\omega^{-2}z-\omega^{-3}z^{2}-\cdots,$
so $(0^{\infty};-\omega^{-1},-\omega^{-2},-\omega^{-3},\cdots)$ becomes the absolutely summable inverse of $z-\omega$.
Case 4: $P(z)=z-w$ with $|\omega|<1$. Since
$(z-\omega)^{-1}=z^{-1}(1-\omega z^{-1})^{-1}=z^{-1}+\omega z^{-2}+\omega^{2}z^{-3}+\cdots,$
then $(\cdots,\omega^{2},\omega,1;0^{\infty})$ becomes the absolutely summable inverse of $z-\omega$.
Thus, $P(z)$ has the absolutely summable inverse which is the convolution product of the above terms. ∎
3 Stream zeros and dynamical systems
Let $P(z)$ be a primitive polynomial. Then clearly $\Omega_{P}$ is a compact $\mathbb{Z}$-subgroup of $St(\mathbb{T})$. Moreover, the pair $(\Omega_{P},\sigma)$, where $\sigma$ is the shift, becomes a compact dynamical system, which is called the dynamical system introduced by $P(z)$. On the other hand, let $\{\Omega_{P}\}=(\{x_{n}\};n\in\mathbb{Z})\in St([0,1))$, where $\{x_{n}\}$ is the representative of $x_{n}\in\mathbb{T}$ in $[0,1)$. Then, $P(z)\times\omega\in St(K_{P})$ holds for any $\omega\in\{\Omega_{P}\}$, where $K_{P}$ is a finite subset of integers with
$K_{P}=\{k_{*},k_{*}+1,\cdots,k^{*}\}$
here
$k_{*}:=(\sum_{n\in S,a_{n}<0}a_{n}+1)\wedge 0\mbox{ and }k^{*}:=(\sum_{n\in S,a_{n}>0}a_{n}-1)\vee 0.$
Let $\Xi_{P}$ be the closure of set $P(z)\times\{\Omega_{P}\}$. Then $\Xi_{P}$ is a compact, shift-invariant subset of $K_{P}^{\mathbb{Z}}$. Together with the shift $\sigma$, we call $(\Xi_{P},\sigma)$ the symbolic representation of $(\Omega_{P},\sigma)$
Theorem 2.
Let a primitive polynomial $P(z)=\sum_{n\in S}a_{n}z^{n}$ be hyperbolic. Then, the dynamical systems $(\Omega_{P},\sigma)$ and $(\Xi_{P},\sigma)$ are conjugate. Hence, we have the following commutative diagram. The conjugate map ×$\times P(z)$ is almost continuous, that is, continuous except for a countable set.
Moreover, a necessary and sufficient condition for $\delta=(\delta(n);~{}n\in\mathbb{Z})\in St(\mathbb{Z})$ to be $\delta\in\Xi_{P}$ is that $P(z)^{-1}\times\delta\in St([0,1))$, which automatically implies that $P(z)^{-1}\times\delta\in\{\Omega_{P}\}$.
$\{\Omega_{P}\}$$\{\Omega_{P}\}$$\Xi_{P}$$\Xi_{P}$$\sigma$$\sigma$×$\times P(z)^{-1}$×$\times P(z)^{-1}$×$\times P(z)$×$\times P(z)$
Proof.
The first half of the statement is clear. We prove the second half. If $\delta\in\Xi_{P}$, then there exists $\omega\in\Omega_{P}$ such that $\delta=P(z)\times(\{\omega_{n}\};n\in\mathbb{Z})$. Therefore, we have
$P(z)^{-1}\times\delta=P(z)^{-1}\times P(z)\times(\{\omega_{n}\};n\in\mathbb{Z})=(\{\omega_{n}\};n\in\mathbb{Z})\in[0,1)^{\mathbb{Z}}.$
Conversely, assume that $\tau:=P(z)^{-1}\times\delta\in[0,1)^{\mathbb{Z}}$. Since $P(z)\times\tau=\delta\in K_{P}^{\mathbb{Z}}$, we have $P(z)\times\pi(\tau)=(0^{\infty};0^{\infty})$, and hence, $\pi(\tau)\in\Omega_{P}$. Here, $\pi:\mathbb{R}^{\infty}\to\mathbb{T}^{\infty}$ is the projection. Since $\tau=\{\pi(\tau)\}$, we have $\delta=P(z)\times\tau\in\Xi_{P}$. ∎
Remark 1.
To recover a dynamical system on $St([0,1))$ from its symbolic representation is known as Kaneko’s formula [4]. It is nothing but to find an absolutely summable $P(z)^{-1}$ as shown in Theorem 2.
Example 1.
Let $P(z)=z^{2}-3z+1$. Then, we have
$\Omega_{P}=\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T}):~{}x_{n+2}=3x_{n+1}-x_{n}~{}\mbox{(mod 1) for any }n\in\mathbb{Z}\}\}.$
We rewrite the above as
$\Omega_{P}=\left\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T}):~{}\left(\begin{array}[]{c}x_{n+1}\\ x_{n+2}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ -1&3\end{array}\right)\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\end{array}\right)\mbox{(mod 1) for any }n\in\mathbb{Z}\right\}.$
Observe that
$P(z)=z^{2}-3z+1=(z-\omega_{+})(z-\omega_{-}),\quad\mbox{here}~{}\omega_{\pm}:=\frac{3\pm\sqrt{5}}{2}.$
Since $|\omega_{+}|=|\frac{3+\sqrt{5}}{2}|>1$,
$(z-\omega_{+})^{-1}=(-1)\omega_{+}^{-1}(1-\omega_{+}^{-1}z)^{-1}=-\omega_{+}^{-1}-\omega_{+}^{-2}z-\omega_{+}^{-3}z^{2}-\cdots,$
Similarly, we can handle the case $|\omega_{-}|=|\frac{3-\sqrt{5}}{2}|<1,$
$(z-\omega_{-})^{-1}=z^{-1}(1-\omega_{-}z^{-1})^{-1}=z^{-1}+\omega_{-}z^{-2}+\omega_{-}^{2}z^{-3}+\cdots,$
together with
$\omega_{+}^{-1}\sum_{n=0}^{\infty}(\omega_{+}^{-1}\omega_{-})^{n}=\omega_{+}^{-1}\frac{1}{1-\omega_{+}^{-1}\omega_{-}}=\frac{1}{\omega_{+}-\omega_{-}}=\frac{1}{\sqrt{5}},$
we have the following Kaneko’s formula:
$P(z)^{-1}=(z-\omega_{+})^{-1}\times(z-\omega_{-})^{-1}$
$=(0^{\infty};-\omega_{+}^{-1},-\omega_{+}^{-2},-\omega_{+}^{-3},\cdots)\times(\cdots,\omega_{-}^{3},\omega_{-}^{2},\omega_{-},1;0^{\infty})$
$=-\frac{1}{\sqrt{5}}(\cdots,\omega_{-}^{3},\omega_{-}^{2},\omega_{-},1;\omega_{+}^{-1},\omega_{+}^{-2},\omega_{+}^{-3},\cdots).$
It is interesting that the above formula can also be obtained by the so-called partial fractional decomposition.
$P(z)^{-1}=(z-\omega_{+})^{-1}\times(z-\omega_{-})^{-1}=\frac{1}{\omega_{+}-\omega_{-}}\{\frac{1}{z-\omega_{-}}-\frac{1}{z-\omega_{-}}\}$
$=-\frac{1}{\sqrt{5}}\left\{\frac{1}{\omega_{+}}\frac{1}{1-z\omega_{+}^{-1}}+\frac{1}{z}\frac{1}{1-z^{-1}\omega_{-}}\right\}$
$=-\frac{1}{\sqrt{5}}\left\{\frac{1}{\omega_{+}}\sum_{n=0}^{\infty}(\omega_{+}^{-1}z)^{n}+\frac{1}{z}\sum_{n=0}^{\infty}(\omega_{-}z^{-1})^{n}\right\}$
$=-\frac{1}{\sqrt{5}}(\cdots,\omega_{-}^{3},\omega_{-}^{2},\omega_{-},1;\omega_{+}^{-1},\omega_{+}^{-2},\omega_{+}^{-3},\cdots).$
$x_{n}$$x_{n+1}$1111011-1- 122-2- 2
Figure 1: Symbolic representation of $St([0,1))$: $\xi_{n+2}\in\{-2,-1,0,1\}.$
[…]
The topological entropy of these dynamical systems coincides with
$\lim_{n\to\infty}\frac{1}{n}\log\#\pi_{n}\Xi_{P},$
where $\pi_{n}$ is the projection $K_{P}^{\mathbb{Z}}\to K_{P}^{\{1,2,\cdots,n\}}$. On the other hand, it is well known that it coincides with the sum of logarithms of the eigenvalues greater than 1 of the companion matrix (see [5], for example). In our case, the companion matrix $\left(\begin{array}[]{cc}0&1\\ -1&3\end{array}\right)$ has eigenvalues $\frac{3\pm\sqrt{5}}{2}$, and hence, the topological entropy is $\log\frac{3+\sqrt{5}}{2}$.
4 Factorization of polynomials and stream zeros
For two primitive polynomials $P(z),Q(z)$ with degree $n$ and $m$, respectively.
$P(z)=a_{0}+a_{1}z+\cdots+a_{n}z^{n}~{}(a_{n}\neq 0)$
$Q(z)=b_{0}+b_{1}z+\cdots+b_{m}z^{m}~{}(b_{m}\neq 0),$
we define the resultant matrix $M$ being a square matrix of degree $n+m$ as follows:
$M=\left(\begin{array}[]{cccccc}a_{0}&a_{1}&\cdots&a_{n}&0&\cdots\\ \cdots&\ddots&\ddots&\ddots&\ddots&\ddots\\ \cdots&0&a_{0}&a_{1}&\cdots&a_{n}\\ b_{0}&b_{1}&\cdots&b_{m}&0&\cdots\\ \cdots&\ddots&\ddots&\ddots&\ddots&\ddots\\ \cdots&0&b_{0}&b_{1}&\cdots&b_{m}\end{array}\right)~{}~{}.$
Then, it is well known [6] that $P(z),Q(z)$ are coprime if and only if $M$ is regular. It also holds that if $P(z),Q(z)$ are coprime, then there exists polynomials $A(z),B(z)$ with integer coefficients such that
$A(z)\times P(z)+B(z)\times Q(z)=\det M\in\mathbb{Z}\setminus\{0\}.$
Denote $\det(M)$ by $Res(P,Q)$. For a stream zero’s of $P(z)$, define $\dim\Omega_{P}$ to be the maximal $k$ such that
$\{(x_{0},x_{1},\cdots,x_{k-1}):~{}(x_{n},n\in\mathbb{Z})\in\Omega_{P}\}=\mathbb{T}^{k}.$
We now show there are close relationships between factorization or irreducibility of a polynomial $P(z)$ and the structure of $\Omega_{P}$.
Theorem 3.
Let $P(z),Q(z)$ be polynomials with integer coefficients.
(i) We have $\dim\Omega_{P}=n-h$, if $P(z)$ has the form $a_{h}z^{h}+a_{h+1}z^{h+1}+\cdots+a_{n}z^{n}$ with $a_{h}\neq 0,a_{n}\neq 0$.
(ii) If $P(z),Q(z)$ are coprime, then we have
$\{\Omega_{P}\}\cap\{\Omega_{Q}\}\subset\{\frac{0}{\Delta},\frac{1}{\Delta},\cdots,\frac{\Delta-1}{\Delta}\}^{\mathbb{Z}}$
with $\Delta=Res(P,Q)$. Therefore, $P(z),Q(z)$ are coprime if and only if $\dim(\Omega_{P}\cap\Omega_{Q})=0$.
(iii) $P(z)$ is a factor of $Q(z)$ if and only if there exists $d\in\mathbb{Z}\setminus\{0\}$ such that $\Omega_{P}\subset\frac{1}{d}\Omega_{Q}$.
(iv) If $P(z),Q(z)$ are coprime, then we have
$\Omega_{P}+\Omega_{Q}\subset\Omega_{P\times Q}\subset\frac{1}{Res(P,Q)}(\Omega_{P}+\Omega_{Q}).$
Hence, we have $\dim\Omega_{P\times Q}=\dim\Omega_{P}+\dim\Omega_{Q}$. Conversely if $P(z),Q(z)$ are not coprime, then we have $\dim\Omega_{P\times Q}>\dim\Omega_{P}+\dim\Omega_{Q}$. Specially if $Res(P,Q)=\pm 1$, then we have $\Omega_{P\times Q}=\Omega_{P}\oplus\Omega_{Q}$, where $\oplus$” is the direct sum.
Proof.
(i) Let $k=n-h$. Since
$(P(z)\times x)_{i}=a_{h}x_{i-h}+a_{h+1}x_{i-h-1}+\cdots+a_{n}x_{i-n}$
for any $(x_{n};n\in\mathbb{Z})\in\mathbb{T}^{\mathbb{Z}}$, a necessary and sufficient condition that for any $(x_{0},x_{n},\cdots,x_{l})\in\mathbb{T}^{\{0,1,\cdots,l\}}$, there exists its extension $(x_{n};n\in\mathbb{Z})\in St(\mathbb{T})$ such that $P(z)\times x=(0^{\infty};0^{\infty})$ is that either $i-h$ or $i-n$ is not in $\{0,1,\cdots,l\}$ for any $i\in\mathbb{Z}$. This is because if so, then we can control the value $x_{i-h}$ or $x_{i-n}$ so that $(P(z)\times x)_{i}=0$. But if not, we have no control. This condition is equivalent to $n-h>l$. Hence, we have $\dim\Omega_{P}=n-h$.
(ii) Let $P(z),Q(z)$ be coprime. Then, the resultant matrix $M$ is regular and the entries of $M^{-1}$ are in $\mathbb{Z}/\Delta$, where $\Delta=Res(P,Q)\in\mathbb{Z}\setminus\{0\}$. Since for any $k\in\mathbb{Z}$ and $\xi=(\xi_{i})_{i\in\mathbb{Z}}\in\{\Omega_{P}\}\cap\{\Omega_{Q}\}$, we have $M\left(\begin{array}[]{c}\xi_{k+1}\\ \xi_{k+2}\\ \cdots\\ \xi_{k+n+m}\end{array}\right)\in\mathbb{Z}^{k+n+m}$, it holds that $\left(\begin{array}[]{c}\xi_{k+1}\\ \xi_{k+2}\\ \cdots\\ \xi_{n+m}\end{array}\right)\in\frac{1}{\Delta}\mathbb{Z}^{n+m}$. This implies that $\{\Omega_{P}\}\cap\{\Omega_{Q}\}\subset\{\frac{0}{\Delta},\frac{1}{\Delta},\cdots,\frac{\Delta-1}{\Delta}\}$, and hence, we have $\dim(\{\Omega_{P}\}\cap\{\Omega_{Q}\})=0$.
Conversely, if $P(z),Q(z)$ are not coprime, then there exists a common factor of them with degree at least 1. Hence we have $\dim(\{\Omega_{P}\}\cap\{\Omega_{Q}\})\geq 1$.
(iii) Assume that $P(z)$ is a factor of $Q(z)$ so that $dQ(z)=P(z)\times R(z)$, where $d\in\mathbb{Z}\setminus\{0\}$ and $R(z)$ is a polynomial with integer coefficients. Then for any $\xi\in\Omega_{P}$, we have
$dQ(z)\times\xi=(P(z)\times R(z))\times\xi=R(z)\times(P(z)\times\xi)=(0^{\infty};0^{\infty}).$
Hence, we have $\xi\in\frac{1}{d}\Omega_{Q}$ and $\Omega_{P}\subset\frac{1}{d}\Omega_{Q}$.
Conversely, if $P(z)$ is not a factor of $Q(z)$, then there exists a nontrivial factor $R(z)$ of $P(z)$ such that $R(z),Q(z)$ are coprime. Then by (ii), there exists $(\xi_{i})_{i\in\mathbb{Z}}\in\Omega_{R}$ such that for any $d\in\mathbb{Z}\setminus\{0\}$, $(\xi_{i})_{i\in\mathbb{Z}}\notin\frac{1}{d}\Omega_{Q}$ holds. Since $(\xi_{i})_{\in\mathbb{Z}}\in\Omega_{P}$, $\Omega_{P}\subset\frac{1}{d}\Omega_{Q}$ does not hold for any $d\in\mathbb{Z}\setminus\{0\}$.
(iv) By (iii) and the fact that $\Omega_{P\times Q}$ is a additive group, it holds that $\Omega_{P\times Q}\supset\Omega_{P}+\Omega_{Q}$. Since $P(z),Q(z)$ are coprime, there exist polynomials with integer coefficients $A(z),B(z)$ such that
$P\times A+Q\times B=\Delta,$
where $\Delta=Res(P,Q).$ Hence, we have
$P\times Q\times B=P\times(\Delta-P\times A),~{}P\times Q\times A=Q\times(\Delta-Q\times B).$
Taking $(x_{n})_{n\in\mathbb{Z}}\in\Omega_{P\times Q}$ arbitrary and by these equalities, we have
$P\times(\Delta-P\times A)\times(x_{n})_{n\in\mathbb{Z}}=Q\times(\Delta-Q\times B)\times(x_{n})_{n\in\mathbb{Z}}=(0^{\infty};0^{\infty}).$
Put
$(u_{n})_{n\in\mathbb{Z}}:=(\Delta-P\times A)\times(x_{n})_{n\in\mathbb{Z}},~{}(v_{n})_{n\in\mathbb{Z}}:=(\Delta-Q\times B)\times(x_{n})_{n\in\mathbb{Z}},$
we have $(u_{n})_{n\in\mathbb{Z}}\in\Omega_{P},~{}(v_{n})_{n\in\mathbb{Z}}\in\Omega_{Q}$ and
$(u_{n})_{n\in\mathbb{Z}}+(v_{n})_{n\in\mathbb{Z}}=(2\Delta-P\times A-Q\times B)\times(x_{n})_{n\in\mathbb{Z}}=\Delta(x_{n})_{n\in\mathbb{Z}}.$
That is, $\Delta(x_{n})_{n\in\mathbb{Z}}\in\Omega_{P}+\Omega_{Q}$, and hence, $\Omega_{P\times Q}\subset\frac{1}{\Delta}(\Omega_{P}+\Omega_{Q})$. Specially if $\Delta=\pm 1$, then we have $\Omega_{P\times Q}=\Omega_{P}+\Omega_{Q}$ and hence, by (ii), $\Omega_{P}\cap\Omega_{Q}=\{0\}$. Thus, $\Omega_{P\times Q}=\Omega_{P}\oplus\Omega_{Q}$ holds.
Conversely, assume that $P(z),Q(z)$ are not coprime. Let $R(z)$ be their GCD and let $P(z)=P^{\prime}(z)R(z),~{}Q(z)=Q^{\prime}(z)R(z)$. Then, we have
$\Omega_{P^{\prime}\times R\times Q^{\prime}}\supset\Omega_{P}+\Omega_{Q}$
since $\Omega_{P^{\prime}\times R\times Q^{\prime}}$ is an additive group containing both $\Omega_{P}$ and $\Omega_{Q}$. Therefore, by (i) and the fact that $P\times Q=P^{\prime}\times R\times Q^{\prime}\times R$ we have
$\dim\Omega_{P\times Q}>\dim\Omega_{P^{\prime}\times R\times Q^{\prime}}\geq\dim(\Omega_{P}+\Omega_{Q}).$
Example 2.
Let us consider the three polynomials $Q(z)=-2z+1,~{}R(z)=-3z+1$ and $P(z)=Q(z)R(z)$. Observe that $Res(P,Q)=1$. Then, we have
$\Omega_{Q}=\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});~{}x_{n+1}=2x_{n}~{}\mbox{(mod 1) for any }n\in\mathbb{Z}\};$
$\Omega_{R}=\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});~{}x_{n+1}=3x_{n}~{}\mbox{(mod 1) for any }n\in\mathbb{Z}\};$
$\Omega_{P}=\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});~{}x_{n+2}=5x_{n+1}-6x_{n}~{}\mbox{(mod 1) for any }n\in\mathbb{Z}\}$
and $\Omega_{Q}\cap\Omega_{R}=(0^{\infty};0^{\infty})$. Moreover, $\Omega_{P}=\Omega_{Q}\oplus\Omega_{R}$ holds.
As a consequence of Theorem 3, we have the following
Corollary 1.
For polynomials $P(z),Q(z),R(z)$ with integer coefficients, assume that $P(z)=Q(z)\times R(z)$ and $Res(Q,R)=\pm 1$. Then dynamical systems $(\Omega_{P},\sigma)$ and $(\Omega_{Q},\sigma)\otimes(\Omega_{R},\sigma)$ are conjugate, where $\otimes$” implies the direct product of dynamical systems. Hence, dynamical systems $(\Xi_{P},\sigma)$ and $(\Xi_{Q},\sigma)\otimes(\Xi_{R},\sigma)$ are conjugate.
5 Stream zeros and the automorphism group
Let $\Omega_{P}$ be the set of stream zero’s of a primitive polynomial $P(z)$. A mapping $\varphi:\Omega_{P}\to\Omega_{P}$ is called an automorphism of $\Omega_{P}$ if it is an isomorphism of $\Omega_{P}$ as the topological additive group such that $\varphi\circ\sigma=\sigma\circ\varphi$. Furthermore, we call an automorphism $\varphi$ a strong automorphism if for any
$(x_{n};n\in\mathbb{Z}),~{}(y_{n};n\in\mathbb{Z})\in\Omega_{P}\mbox{ with }\varphi(x_{n};n\in\mathbb{Z})=(y_{n};n\in\mathbb{Z}),$
$(y_{n};n\in\mathbb{Z}_{\geq 0})$ is determined by $(x_{n};n\in\mathbb{Z}_{\geq 0})$ and $(x_{n};n\in\mathbb{Z}_{\geq 0})$ is determined by $(y_{n};n\in\mathbb{Z}_{\geq 0})$. Clearly, the shift $\sigma$ is an automorphism and $I,-I$ are strong automorphisms, where ±$\pm I$ are the convolution products by ±$\pm I$.
Let $P(z)=a_{k}z^{k}+a_{k-1}z^{k-1}+\cdots+a_{1}z+1\in\mathbb{Z}[z]$ with $k\geq 1,a_{k}\neq 0$. That is,
$P(z)=(0^{\infty};1,a_{1},a_{2},\cdots,a_{k-1},a_{k},0^{\infty})\in St(\mathbb{Z}).$ (3)
Therefore,
$\Omega_{P}=\{(x_{n};n\in\mathbb{Z})\in\mathbb{T}^{\mathbb{Z}}:~{}x_{n+k}=-a_{1}x_{n+k-1}-a_{2}x_{n+k-2}-\cdots-a_{k}x_{n}~{}(\forall n\in\mathbb{Z})\}.$
Then, we rewrite the above as
$\Omega_{P}=\left\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T}):~{}\left(\begin{array}[]{c}x_{n+1}\\ x_{n+2}\\ \cdot\\ \cdot\\ x_{n+k}\end{array}\right)=M_{P}\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\\ \cdot\\ \cdot\\ x_{n+k-1}\end{array}\right)\mbox{ for any }n\in\mathbb{Z}\right\}.$
where $M_{P}$ denote the associate matrix as follows:
$M_{P}:=\left(\begin{array}[]{ccccc}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \cdots&\cdots&\cdots&\ddots&\cdots\\ 0&0&0&\cdots&1\\ -a_{k}&-a_{k-1}&-a_{k-2}&\cdots&-a_{1}\end{array}\right).$
Note that $M_{P}$ is not the companion matrix of the polynomial $P(z)$. Indeed, it is the companion matrix of the reciprocal polynomial of $P(z)$.
Definition 1.
The set of $k\times k$ integer matrices B such that $BM_{P}=M_{P}B$ and $\det B=\pm 1$ is denoted by $Saut_{P}$.
Recall that a square $k\times k$ integer matrix having determinant ±$\pm 1$ is called unimodular matrix. Equivalently, it is an integer matrix that is invertible over the integers. In group theory, the $k\times k$ unimodular matrices under matrix multiplication form a group called the $k\times k$ general linear group over $\mathbb{Z}$, which is denoted $GL(k,\mathbb{Z})$. Therefore,
$Saut_{P}=\{B\in GL(k,\mathbb{Z}):BM_{P}=M_{P}B\}.$
Observe that $Saut_{P}$ is a subgroup of $GL(k,\mathbb{Z})$. In fact, when $BM_{P}=M_{P}B$, then
$B^{-1}M_{P}=B^{-1}M_{P}BB^{-1}=B^{-1}BM_{P}B^{-1}=M_{P}B^{-1}.$
Furthermore, if $B_{1}M_{P}=M_{P}B_{1}$ and $B_{2}M_{P}=M_{P}B_{2}$, then
$B_{1}B_{2}M_{P}=B_{1}M_{P}B_{2}=M_{p}B_{1}B_{2}.$
Note that the $k\times k$ identity matrix $I_{k}$ always be in $Saut_{P}$, while the matrix $M_{p}$ may not be in $GL(k,\mathbb{Z})$, if $M_{p}\in GL(k,\mathbb{Z})$, then $Saut_{P}$ contains all powers of $M_{p}$ and their reflections. This happens to constitute all the elements of $Saut_{P}$ in some special cases, see the Example 4 below.
The following theorem states that each of strong automorphisms on $\Omega_{P}$ can be represented by a matrix from $Saut_{P}$.
Theorem 4.
Assume $(3)$. Then $\varphi:\Omega_{P}\to\Omega_{P}$ is a strong automorphism if and only if there exists a $k\times k$ integer matrix $B\in Saut_{P}$ such that
$\left(\begin{array}[]{c}y_{n+1}\\ y_{n+2}\\ \cdot\\ \cdot\\ y_{n+k}\end{array}\right)=B\left(\begin{array}[]{c}x_{n+1}\\ x_{n+2}\\ \cdot\\ \cdot\\ x_{n+k}\end{array}\right)\mbox{for any $n\in\mathbb{Z}$,}$ (4)
and $(x_{n};n\in\mathbb{Z}),(y_{n};n\in\mathbb{Z})\in\Omega_{P}$ with $\varphi(x_{n};n\in\mathbb{Z})=(y_{n};n\in\mathbb{Z})$.
Proof.
Let $B\in Saut_{P}$ and satisfying (4). By the hypothesis of $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$, we have
$\left(\begin{array}[]{c}y_{n+1}\\ y_{n+2}\\ \cdot\\ \cdot\\ y_{n+k}\end{array}\right)=B\left(\begin{array}[]{c}x_{n+1}\\ x_{n+2}\\ \cdot\\ \cdot\\ x_{n+k}\end{array}\right)=BM_{P}\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\\ \cdot\\ \cdot\\ x_{n+k-1}\end{array}\right)\\ =M_{P}B\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\\ \cdot\\ \cdot\\ x_{n+k-1}\end{array}\right)=M_{P}\left(\begin{array}[]{c}y_{n}\\ y_{n+1}\\ \cdot\\ \cdot\\ y_{n+k-1}\end{array}\right),$
thus $(y_{n};n\in\mathbb{Z})\in\Omega_{P}$ is well defined. On the other hand, we already know that $B^{-1}$ exists and $B^{-1}M_{P}=M_{P}B^{-1}.$ So this correspondence is clearly a strong automorphism which will be denoted as $\varphi_{B}$.
Conversely, if $\varphi$ is a strong automorphism and $(y_{n};n\in\mathbb{Z})=\varphi(x_{n};n\in\mathbb{Z})$, then $(x_{i};n\leq i\leq n+k-1)$ determines $(x_{i};i\geq n)$ which determines $(y_{i};i\geq n)$, and hence $(y_{i};n\leq i\leq n+k-1)$. This converse also holds, and the correspondence between $(x_{i};n\leq i\leq n+k-1)\in\mathbb{T}^{k}$ and $(y_{i};n\leq i\leq n+k-1)\in\mathbb{T}^{k}$ is automorphism, it is described as $\varphi_{B}$ with $B\in Saut_{P}$. ∎
Due to Theorem 4, we turn to investigate the structure of $Saut_{P}$.
Theorem 5.
Assume that all roots $\theta_{1},\cdots,\theta_{k}$ of $P(z)=a_{k}z^{k}+a_{k-1}z^{k-1}+\cdots+a_{1}z+1\in\mathbb{Z}[z],k\geq 1,a_{k}\neq 0$ are simple, then we have the following:
(i) For a $k\times k$ matrix $B$, $B\in Saut_{P}$ if and only if there exist $c_{i}\in\mathbb{Z}~{}(i=0,1,\cdots,k-1)$ such that for $\lambda_{i}=c_{0}+c_{1}\theta_{i}^{-1}+\cdots+c_{k-1}\theta_{i}^{-(k-1)}~{}(i=1,\cdots,k)$ with $\lambda_{1}\lambda_{2}\cdots\lambda_{k}=\pm 1$,
$B=\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)\left(\begin{array}[]{ccc}\lambda_{1}&0&\cdots\\ 0&\ddots&0\\ \cdots&0&\lambda_{k}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)^{-1}$ (5)
holds. Hence, $Saut_{P}$ is an abelian group.
(ii) Assume additionally that all the roots of $P(z)=0$ are real, then for any $B\in Saut_{P}$, the order of $B$ is either $1,2$ or infinity. The set of $B\in Saut_{P}$ with finite order is a finite subgroup ${\mathcal{F}}_{P}$ of $Saut_{P}$, and the quotient group $Saut_{P}/{\mathcal{F}}_{P}$ is either trivial or a countable multiplicative subroup of $\mathbb{R}_{>0}$.
Proof.
(i) Note that $\theta_{i}\neq 0~{}(i=1,2,\cdots,k)$ by (3). Assume $B\in Saut_{P}$. Then it is an integer matrix. Since
$M_{P}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=\left(\begin{array}[]{c}\theta_{i}^{k-1}\\ \vdots\\ \theta_{i}\\ 1\end{array}\right)=\theta_{i}^{-1}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)~{}(i=1,2,\cdots,k)$
and $BM_{P}=M_{P}B$, we have
$M_{P}B\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=BM_{P}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=\theta_{i}^{-1}B\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)~{}(i=1,2,\cdots,k).$
Since each eigenspace of $M_{P}$ is 1-dimensional, there exists $\lambda_{i}\in\mathbb{C}$ such that
$B\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=\lambda_{i}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)~{}(i=1,2,\cdots,k).$
This implies (5). Moreover, let $(c_{0},c_{1},\cdots,c_{k-1})$ be the first row of $B$, then from the above, we have $c_{0}\theta_{i}^{k}+c_{1}\theta_{i}^{k-1}+\cdots+c_{k-1}\theta_{i}=\lambda_{i}\theta_{i}^{k}$, so that
$\lambda_{i}=c_{0}+c_{1}\theta_{i}^{-1}+\cdots+c_{k-1}\theta_{i}^{-(k-1)}~{}(i=1,\cdots,k).$ (6)
Conversely, if $B$ is a $k\times k$ matrix satisfying (5) and (6) with $\lambda_{i}\in\mathbb{C}~{}(i=1,\cdots,k)$ such that $\lambda_{1}\cdots\lambda_{k}=\pm 1$. Then, we have $\det B=\pm 1$. Moreover, since
$M_{P}=\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{-1}&0&\cdots\\ 0&\ddots&0\\ \cdots&0&\theta_{k}^{-1}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)^{-1},$
we have $BM_{P}=M_{P}B,$ since both terms are equal to
$\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)\left(\begin{array}[]{ccc}\lambda_{1}\theta_{1}^{-1}&0&\cdots\\ 0&\ddots&0\\ \cdots&0&\lambda_{k}\theta_{k}^{-1}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)^{-1}.$
Thus, we have $B\in Saut_{P}$. Note that all entries of $B$ are integers since $B$ is a linear combination of
$M_{P}^{l}=\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{-l}&0&\cdots\\ 0&\ddots&0\\ \cdots&0&\theta_{k}^{-l}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)^{-1},$
with integer coefficients $c_{0},c_{1},\cdots,c_{k-1}$ by (6).
(ii) Assume that all the roots of $P(z)=0$ are real. Then, (5) gives an isomorphism from $Saut_{P}$ into
$\{(\lambda_{1},\cdots,\lambda_{k})\in(\mathbb{R}\setminus\{0\})^{k};~{}\lambda_{1}\cdots\lambda_{k}=\pm 1\}.$
If $B$ is mapped to one of them with $\lambda_{i}\neq\pm 1$ for some $i=1,2,\cdots,k$, then it has an infinite order. Otherwise, it is mapped to $(\lambda_{1},\cdots,\lambda_{k})$ with $\lambda_{i}\in\{\pm 1\}~{}(i=1,\cdots,k)$. Hence, the order is either $1$ or $2$.
The final statement follows since $\lambda_{1}$ in (5) determine $B$ and the set of $\lambda_{1}$ of the form $\lambda_{1}=c_{0}+c_{1}\theta_{1}^{-1}+\cdots+c_{k-1}\theta_{1}^{-(k-1)}\notin\{0,\pm 1\}$ with $c_{i}\in\mathbb{Z}~{}(i=0,1,\cdots,k-1)$ is a countable multiplicative subgroup of $\mathbb{R}_{>0}$. ∎
6 $Saut_{P}$ in the case $\deg P=2,3$
We now give a complete characterization of the elements of $Saut_{P}$ when the degree of polynomial $P(z)$ is two. More precisely, we have
Theorem 6.
Assume $(3)$ with the degree $k=2$ and discriminant $D=a_{1}^{2}-4a_{2}$.
(i) Assume $D>0$. Then the following conditions are equivalent:
(i-1) $B\in Saut_{P}$,
(i-2) $B$ is a $2\times 2$ integer matrix such that $(5)$ holds for some $\lambda_{i}\in\mathbb{C}~{}(i=1,2)$ such that $\lambda_{1}\lambda_{2}=1$,
(i-3)  $B=\left(\begin{array}[]{cc}c_{0}&c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right)$ for some $c_{0},c_{1}\in\mathbb{Z}$ such that
$(2c_{0}-a_{1}c_{1})^{2}-Dc_{1}^{2}=\pm 4,$ (7)
(i-4) $B$ is of the form $\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)\in GL(2,\mathbb{Z})$ satisfying
$\frac{p\theta+p^{\prime}}{q\theta+q^{\prime}}=\theta$ (8)
for some root $\theta$ of the equation $a_{2}z^{2}+a_{1}z+1=0$.
Moreover, $Saut_{P}/\{\pm I_{2}\}$ is an infinite cyclic group if and only if $D$ is a non-square integer. In addition, if $D=4$, then it is a cyclic group of order $2$. Otherwise, it is the trivial group.
(ii) If $D=0$, then $Saut_{P}/\{\pm I_{2}\}$ is an infinite cyclic group.
(iii) Assume $D<0$. Then $Saut_{P}$ is the trivial group except for the case $D=-4$, where $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order $2$.
Remark 2.
The essential part of (ii) of Theorem 6 is to solve the so-called Pell’s equation, i.e., Diophantine equation of the form
$w^{2}-Dv^{2}=\pm 4,\ w,v\in Z,~{}\mbox{for a given integer}\,D.$ (9)
It is well known that the main tool to solve it is the theory of continued fraction (see [1, 3], for example).
Proof.
(i) Assume $D>0$. Let $\theta_{i}~{}(i=1,2)$ be the two roots of $P(z)=0$. By Theorem 5, $B=\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)\in Saut_{p}$ if and only if $pq^{\prime}-p^{\prime}q=\pm 1$ and that
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}\theta_{1}^{2}&\theta_{2}^{2}\\ \theta_{1}&\theta_{2}\end{array}\right)\left(\begin{array}[]{cc}\lambda_{1}&0\\ 0&\lambda_{2}\end{array}\right)\left(\begin{array}[]{cc}\theta_{1}^{2}&\theta_{2}^{2}\\ \theta_{1}&\theta_{2}\end{array}\right)^{-1}$ (10)
holds for some $\lambda_{i}~{}(i=1,2)$ such that $\lambda_{i}=c_{0}+c_{1}\theta_{i}^{-1}$ with $c_{0},c_{1}\in\mathbb{Z}$. Combing with $\theta_{1}\theta_{2}=\frac{1}{a_{2}},~{}\theta_{1}+\theta_{2}=-\frac{a_{1}}{a_{2}}$, we thus have
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}c_{0}&c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right).$
Furthermore, $pq^{\prime}-p^{\prime}q=\pm 1$ is equivalent to (7). Thus, (i-1) and (i-3) are equivalent. Here, the additional condition that $\lambda_{i}=c_{0}+c_{1}\theta_{i}^{-1}$ with $c_{0},c_{1}\in\mathbb{Z}$ follows immediately if we assume that $B$ is an integer matrix satisfying (10). Hence, they are also equivalent to (i-2).
Let $\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)\in Saut_{p}$. Then by (10), we have
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)\left(\begin{array}[]{cc}\theta_{1}^{2}&\theta_{2}^{2}\\ \theta_{1}&\theta_{2}\end{array}\right)=\left(\begin{array}[]{cc}\theta_{1}^{2}&\theta_{2}^{2}\\ \theta_{1}&\theta_{2}\end{array}\right)\left(\begin{array}[]{cc}\lambda_{1}&0\\ 0&\lambda_{2}\end{array}\right).$
for some $\lambda_{i}~{}(i=1,2)$ with $\lambda_{1}\lambda_{2}=\pm 1$. It follows that
$\frac{p\theta_{i}+p^{\prime}}{q\theta_{i}+q^{\prime}}=\theta_{i}\mbox{ and }pq^{\prime}-p^{\prime}q=\pm 1~{}~{}(i=1,2).$
Thus, (i-1) implies (i-4).
Conversely, assume (i-4). Without loss of generality, assume $\frac{p\theta_{1}+p^{\prime}}{q\theta_{1}+q^{\prime}}=\theta_{1}$. Then since $\theta_{1}$ and $\theta_{2}$ are algebraically conjugate, we also have $\frac{p\theta_{2}+p^{\prime}}{q\theta_{2}+q^{\prime}}=\theta_{2}$. Therefore, there exists $\lambda_{i}~{}(i=1,2)$ such that (10) holds and (i-4) implies (i-2). Thus, the first half of statement (i) is proved.
Let $D>0$. Consider the equation (7), it always has trivial solutions $c_{0}=\pm 1,~{}c_{1}=0$, that correspond to $B=\pm I_{2}$. If $D$ is a square integer, then it is easy to see that (7) has a nontrivial solution only if $D=4$, since the difference of 2 square integers is 4 only if they are 0 and 4. Assume that $Saut_{P}/\{\pm I_{2}\}$ has an element, say $B$ of order $2$. This implies that
$B=\left(\begin{array}[]{cc}\theta_{1}^{2}&\theta_{2}^{2}\\ \theta_{1}&\theta_{2}\end{array}\right)\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)\left(\begin{array}[]{cc}\theta_{1}^{2}&\theta_{2}^{2}\\ \theta_{1}&\theta_{2}\end{array}\right)^{-1}$
is an integer matrix. Hence,
$M_{P}B+M_{P}=\frac{1}{\theta_{1}\theta_{2}(\theta_{1}-\theta_{2})}\left(\begin{array}[]{cc}2\theta_{1}\theta_{2}&-2\theta_{1}\theta_{2}^{2}\\ -2\theta_{2}&-2\theta_{2}^{2}\end{array}\right)$
is an integer matrix. Therefore, $\theta_{1}$ and $\theta_{2}$ are rational since they are ratios of its entries. By $\theta_{1},\theta_{2}=\frac{-a_{1}\pm\sqrt{D}}{2}$, this implies that $D$ is a square integer.
Thus, if $D$ is a nonsquare integer, then $Saut_{P}/\{\pm I_{2}\}$ except for the unit contains only elements of infinite order, and later we prove that it does contain elements of infinite order. Let $\eta$ be the mapping defined on $Saut_{P}$ such that for $B\in Saut_{P}$, $\eta(B)=(\lambda_{1},\lambda_{2})$, where $(\lambda_{1},\lambda_{2})$ satisfies (6) with $k=2$. Then $\eta(B)\in(\mathbb{R}\setminus\{0\})^{2}$ and $Image(\eta)$ is a discrete multiplicative subgroup of ($(\mathbb{R}\setminus\{0\})^{2}$. Let $\tilde{\lambda}$ be the smallest element in $\{\lambda_{1}\in(1,\infty);~{}(\lambda_{1},\lambda_{2})\in Image(\eta)\}$. Since $(1,-1)\notin\eta(B)$, just one of $(\tilde{\lambda},\tilde{\lambda}^{-1})$ or $(\tilde{\lambda},-\tilde{\lambda}^{-1})$ is in $Image(\eta)$. This element generates $\tilde{\lambda}$.
Let $D>0$ be a nonsquare integer. We prove that $Saut_{P}/\{\pm I_{2}\}$ contains nontrivial elements using the continued expansion of $\theta$, which is ultimately periodic. Let it be
$\theta_{1}=[c_{0};c_{1},\cdots,c_{k-1},(e_{0},e_{1}\cdots,e_{m-1})^{\infty}].$
Then,
$\frac{C}{G}=[c_{0};c_{1},\cdots,c_{k-1}],~{}\frac{C^{\prime}}{G^{\prime}}=[c_{0};c_{1},\cdots,c_{k-2}]$
$\frac{E}{F}=[e_{0};e_{1},\cdots,e_{m-1}],~{}\frac{E^{\prime}}{F^{\prime}}=[e_{0};e_{1},\cdots,e_{m-2}],$
be the irreducible fractions with nonnegative denominators, where if $k=0$ and the expansion is purely periodic, then define $\frac{C}{G}=\frac{1}{0},~{}\frac{C^{\prime}}{G^{\prime}}=\frac{0}{1}$, and if $k=1$ ($m=1$), then define $\frac{C^{\prime}}{G^{\prime}}=\frac{1}{0}$ ($\frac{E^{\prime}}{F^{\prime}}=\frac{1}{0}$). Finally, define $2\times 2$-matrix $\left(\begin{array}[]{cc}p_{n}&p^{\prime}_{n}\\ q_{n}&q^{\prime}_{n}\end{array}\right)~{}(n=0,1,2\cdots)$ by
$\left(\begin{array}[]{cc}p_{n}&p^{\prime}_{n}\\ q_{n}&q^{\prime}_{n}\end{array}\right)=\left(\begin{array}[]{cc}C&C^{\prime}\\ G&G^{\prime}\end{array}\right)\left(\begin{array}[]{cc}E&E^{\prime}\\ F&F^{\prime}\end{array}\right)^{n}\left(\begin{array}[]{cc}C&C^{\prime}\\ G&G^{\prime}\end{array}\right)^{-1}.$
Then, they are elements in $Saut_{P}$ different from each other for $n=0,1,\cdots$ so that $Saut_{P}/\{\pm I_{2}\}$ is an infinite cyclic group.
Let $D=4$. Since $D=a_{1}^{2}-4a_{2}$, $a_{1}$ is even, say $a_{1}=2a,a\in\mathbb{Z}$. Then by (10), $(p-ap^{\prime})^{2}-p^{\prime 2}=\pm 1$. Hence, either $p^{\prime}=\pm 1,~{}p=\pm a$ ($\pm$ correspondingly) or $p^{\prime}=0,~{}p=\pm 1$. In the latter case, we have $B=\pm I_{2}$. In the former case, we have $a_{2}=a^{2}-1$, because $D/4=a^{2}-a_{2}=1$. Then by (9),
$B=\pm\left(\begin{array}[]{cc}a&1\\ -a^{2}+1&-a\end{array}\right).$
Note that $\det B=-1$ and $B^{2}=I_{2}$. Thus, $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order 2.
Let $D>0$ be a square other than 4. By (11), $B\in Saut_{P}$ if and only if $(2p-a_{1}p^{\prime})^{2}-Dp^{\prime 4}=\pm 4$, which is possible only $p=1,~{}p^{\prime}=0$. Thus, $Saut_{P}/\{\pm I_{2}\}$is trivial.
(ii) Let $D=0$. Then by (10), $B\in Saut_{P}$ if and only if
$B=\left(\begin{array}[]{cc}p&p^{\prime}\\ -a_{2}p^{\prime}&p-a_{1}p^{\prime}\end{array}\right)\mbox{ with }\det B=\left(p-\frac{a_{1}p^{\prime}}{2}\right)^{2}=1.$
Since $D=a_{1}^{2}-4a_{2}=0$, $a_{1}$ is even. Put $a_{1}=2c$. Then, $p-cp^{\prime}=\pm 1$ and
$B=\left(\begin{array}[]{cc}\pm 1+cp^{\prime}&p^{\prime}\\ -c^{2}p^{\prime}&\pm 1-cp^{\prime}\end{array}\right)\in Saut_{P}~{}(\forall p^{\prime}\in\mathbb{Z})~{}(\mbox{$\pm$ correspondingly}).$
Thus, we have Case 3 and $Saut_{P}/\{\pm I_{2}\}$ is an infinite cyclic group.
(iii) Note that $D\equiv 1~{}(\mbox{mod}~{}4)$. Hence, $D\leq-3$. Then by (7), we should have $c_{1}=0$ or $c_{1}=\pm 1$. If $c_{1}=0$, then $c_{0}=\pm 1$ and hence, $B=\pm I_{2}$. If $c_{1}=\pm 1$, then $D=-4$. Therefore, $Saut_{P}$ is trivial except for the case $D=-4$. If $D=-4$, $Saut_{P}$ contains other elements comming from the case $c=\pm 1$. In fact, in this case, we have $c_{1}=\pm 1,~{}c_{0}=\pm\frac{a_{1}}{2}$ ($\pm$ correspondingly). Note that since $D=a_{1}^{2}-4a_{2}=-4$, $a_{1}$ is even. Hence, $Saut_{P}$ consists of ±$\pm I_{2}$ and ±$\pm\left(\begin{array}[]{cc}\frac{1}{2}a_{1}&1\\ -a_{2}&-\frac{1}{2}a_{1}\end{array}\right)$. Since $\left(\begin{array}[]{cc}\frac{1}{2}a_{1}&1\\ -a_{2}&-\frac{1}{2}a_{1}\end{array}\right)^{2}=I_{2}$, $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order 2. ∎
We now give several examples to illustrate Theorem 6.
Example 3.
In the case of $D>0$, there are three sub-cases: i) $D$ non-square; ii) $D$ square, D = 4; iii) $D$ square, $D\neq 4$.
For i), Let $P(z)=z^{2}-3z+1\in\mathbb{Z}[z]$, then $D=5>0$ and non-square. It is clear that $\theta=\frac{\sqrt{5}+3}{2}$ satisfying $\theta^{2}-3\theta+1=0$, and has continued fraction expansion $\theta=[2;1,1,\cdots]$. Then, we have
$\frac{C}{G}=\frac{2}{1},~{}\frac{C^{\prime}}{G^{\prime}}=\frac{1}{0},~{}\frac{E}{F}=\frac{1}{1},~{}\frac{E^{\prime}}{F^{\prime}}=\frac{1}{0}$
and
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}2&1\\ 1&0\end{array}\right)\left(\begin{array}[]{cc}1&1\\ 1&0\end{array}\right)\left(\begin{array}[]{cc}2&1\\ 1&0\end{array}\right)^{-1}=\left(\begin{array}[]{cc}2&-1\\ 1&-1\end{array}\right).$
Noting that $a_{2}=-1,~{}a_{1}=3$ and the solution of (9) is $v=w=1$, the solution of (7) is $B=\left(\begin{array}[]{cc}-1&1\\ -1&2\end{array}\right)$ and $Saut_{P}=\{\pm B^{n};n\in\mathbb{Z}\}$.
Let $X=(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $\varphi_{B}(X)=(y_{n};n\in\mathbb{Z})$. Then,
$\left(\begin{array}[]{c}y_{0}\\ y_{1}\end{array}\right)=\left(\begin{array}[]{cc}-1&1\\ -1&2\end{array}\right)\left(\begin{array}[]{c}x_{0}\\ x_{1}\end{array}\right)=\left(\begin{array}[]{c}-x_{0}+x_{1}\\ -x_{0}+2x_{1}\end{array}\right).$
Since $y_{2}=-y_{0}+3y_{1}=-2x_{0}+5x_{1}$, we have
$\left(\begin{array}[]{c}y_{1}\\ y_{2}\end{array}\right)=\left(\begin{array}[]{cc}-1&1\\ -1&2\end{array}\right)\left(\begin{array}[]{c}x_{1}\\ -x_{0}+3x_{1}\end{array}\right)=\left(\begin{array}[]{c}-x_{0}+2x_{1}\\ -2x_{0}+5x_{1}\end{array}\right)$
as expected.
For ii), let $P(z)=8z^{2}+6z+1\in\mathbb{Z}[z]$, then $D=4>0.$ $B=\left(\begin{array}[]{cc}3&1\\ -8&-3\end{array}\right).$ Clearly, we have $\det B=-1$ and $B^{2}=I_{2}$. So $Saut_{P}/\{\pm I_{2}\}$ is an finite cyclic group of order 2.
For iii), let $P(z)=-4z^{2}+3z+1\in\mathbb{Z}[z]$, then $D=25$, which is a square and not equal to 4. $B=\pm I_{2}$, so the group $Saut_{P}/\{\pm I_{2}\}$ is trivial.
Example 4.
Let $P(z)=-3z^{2}+1$ with $D=12$. Then $\theta_{1}=\frac{1}{\sqrt{3}}$ is a root of $P(z)=0$ having the continued fraction expansion $\theta_{1}=[0;1,1,2,1,1,2,\cdots]$. Then, we have
$\frac{C}{G}=\frac{0}{1},~{}\frac{C^{\prime}}{G^{\prime}}=\frac{1}{0},~{}\frac{E}{F}=\frac{5}{3},~{}\frac{E^{\prime}}{F^{\prime}}=\frac{1}{0}$
and
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\left(\begin{array}[]{cc}5&1\\ 0&3\end{array}\right)\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)^{-1}=\left(\begin{array}[]{cc}2&1\\ 3&2\end{array}\right).$
Noting that $a_{2}=3,~{}a_{1}=0$ and $v=1,~{}w=4$ in (9), the solution of (8) is obtained as $B=\left(\begin{array}[]{cc}2&-1\\ -3&2\end{array}\right)$ and $Saut_{P}=\{\pm B^{n};n\in\mathbb{Z}\}$.
Let $X=(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $\varphi_{B}(X)=(y_{n};n\in\mathbb{Z})$. Then,
$\left(\begin{array}[]{c}y_{0}\\ y_{1}\end{array}\right)=\left(\begin{array}[]{cc}2&-1\\ -3&2\end{array}\right)\left(\begin{array}[]{c}x_{0}\\ x_{1}\end{array}\right)=\left(\begin{array}[]{c}2x_{0}-x_{1}\\ -3x_{0}+2x_{1}\end{array}\right).$
Since $y_{2}=3y_{0}=6x_{0}-3x_{1}$, we have
$\left(\begin{array}[]{c}y_{1}\\ y_{2}\end{array}\right)=\left(\begin{array}[]{cc}2&-1\\ -3&2\end{array}\right)\left(\begin{array}[]{c}x_{1}\\ 3x_{0}\end{array}\right)=\left(\begin{array}[]{c}-3x_{0}+2x_{1}\\ 6x_{0}-3x_{1}\end{array}\right)$
as expected.
Example 5.
Let $P(z)=z^{2}-2z+1$ with $D=0$. Then by Theorem 5, $Saut_{P}=\{\pm B^{n};n\in\mathbb{Z}\}$ with $B=\left(\begin{array}[]{cc}0&1\\ -1&2\end{array}\right)$. Let $X=(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $\varphi_{B}(X)=(y_{n};n\in\mathbb{Z})$. Then,
$\left(\begin{array}[]{c}y_{0}\\ y_{1}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ -1&2\end{array}\right)\left(\begin{array}[]{c}x_{0}\\ x_{1}\end{array}\right)=\left(\begin{array}[]{c}x_{1}\\ -x_{0}+2x_{1}\end{array}\right).$
Since $x_{2}=-x_{0}+2x_{1},~{}y_{2}=-y_{0}+2y_{1}=-2x_{0}+3x_{1}$, we have
$\left(\begin{array}[]{c}y_{1}\\ y_{2}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ -1&2\end{array}\right)\left(\begin{array}[]{c}x_{1}\\ -x_{0}+2x_{1}\end{array}\right)=\left(\begin{array}[]{c}-x_{0}+2x_{1}\\ -2x_{0}+3x_{1}\end{array}\right)$
as expected.
Example 6.
In the case of $D<0$, there are two sub-cases:
$(i)~{}D=-4\mbox{ and }(ii)~{}D\neq-4.$
For (i), take $P(z)=z^{2}+1\in\mathbb{Z}[z]$, then $D=-4<0.$ It is easy to show that $B=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right).$ we have $\det B=1$ and $B^{2}=I_{2}$. So $Saut_{P}/\{\pm I_{2}\}$ is an finite cyclic group of order 2.
For (ii), choose $P(z)=z^{2}+3z+3\in\mathbb{Z}[z]$, then $D=-3<0.$ As before, $B=\pm I_{2}.$ So $Saut_{P}/\{\pm I_{2}\}$ is an trivial group.
We now discuss the case of degree 3. It’s much harder than the case of degree 2 and we have a result only for a special case.
Theorem 7.
(i) Assume that the roots $\theta_{1},\theta_{2},\theta_{3}$ of $P(z)=a_{3}z^{3}+a_{2}z^{2}+a_{1}z+1\in\mathbb{Z}[z],a_{3}\neq 0$ are simple, then $B\in Saut_{P}$ if and only if
$B=\left(\begin{array}[]{ccc}c_{0}&c_{1}&c_{2}\\ -a_{3}c_{2}&c_{0}-a_{2}c_{2}&c_{1}-a_{1}c_{2}\\ -a_{3}c_{1}+a_{1}a_{3}c_{2}&-a_{2}c_{1}+(a_{1}a_{2}-a_{3})c_{2}&c_{0}-a_{1}c_{1}+(a_{1}^{2}-a_{2})c_{2}\end{array}\right)$
for some $c_{0},c_{1},c_{2}\in\mathbb{Z}$ such that
$\begin{split}~{}~{}~{}\det B&=c_{0}^{3}-a_{3}c_{1}^{3}+a_{3}^{2}c_{2}^{3}-a_{1}c_{0}^{2}c_{1}+(a_{1}^{2}-2a_{2})c_{0}^{2}c_{2}+a_{2}c_{0}c_{1}^{2}\\ &+(a_{2}^{2}-2a_{1}a_{3})c_{0}c_{2}^{2}+a_{1}a_{3}c_{1}^{2}c_{2}-a_{2}a_{3}c_{1}c_{2}^{2}+(3a_{3}-a_{1}a_{2})c_{0}c_{1}c_{2}\\ &=\pm 1.\end{split}$
(ii) Let $P(z)=-az^{3}+1\in\mathbb{Z}[z]$. Then, $B\in Saut_{P}$ if and only if $B=\left(\begin{array}[]{ccc}c_{0}&c_{1}&c_{2}\\ ac_{2}&c_{0}&c_{1}\\ ac_{1}&ac_{2}&c_{0}\end{array}\right)$ for some $c_{0},c_{1},c_{2}\in\mathbb{Z}$ such that
$\det B=c_{0}^{3}+ac_{1}^{3}+a^{2}c_{2}^{3}-3ac_{0}c_{1}c_{2}=\pm 1.$ (11)
(iii) Let $a=1$ in (i). Then, $Saut_{P}/\{\pm I_{3}\}$ is a cyclic group of order $3$ spanned by $\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)$.
Proof.
(i) By hypothesis, $\theta_{i}^{-1}(i=1,2,3)$ is the roots of the reciprocal polynomial of $p(z)$, say $h(z)$, i.e., $h(z)=z^{3}+a_{1}z^{2}+a_{2}z+a_{3}$. Put $\tau_{1}:=\frac{1}{\theta_{1}}+\frac{1}{\theta_{2}}+\frac{1}{\theta_{3}}$, $\tau_{2}:=\frac{1}{\theta_{1}\theta_{2}}+\frac{1}{\theta_{1}\theta_{3}}+\frac{1}{\theta_{2}\theta_{3}}$, $\tau_{3}:=\frac{1}{\theta_{1}\theta_{2}\theta_{3}}$.
Applying theorem 5 with $k=3$,
$B=\left(\begin{array}[]{ccc}\theta_{1}^{3}&\theta_{2}^{3}&\theta_{3}^{3}\\ \theta_{1}^{2}&\theta_{2}^{2}&\theta_{3}^{2}\\ \theta_{1}&\theta_{2}&\theta_{3}\end{array}\right)\left(\begin{array}[]{ccc}\lambda_{1}&0&0\\ 0&\lambda_{2}&0\\ 0&0&\lambda_{3}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{3}&\theta_{2}^{3}&\theta_{3}^{3}\\ \theta_{1}^{2}&\theta_{2}^{2}&\theta_{3}^{2}\\ \theta_{1}&\theta_{2}&\theta_{3}\end{array}\right)^{-1}$
where $\lambda_{i}=c_{0}+c_{1}\theta_{i}^{-1}+c_{2}\theta_{i}^{-2}~{}(i=1,2,3).$ A careful calculation from above gives
$B=\left(\begin{array}[]{ccc}c_{1}&c_{2}&c_{3}\\ c_{2}\tau_{3}&c_{0}-c_{2}\tau_{2}&c_{1}+c_{2}\tau_{1}\\ c_{1}\tau_{3}+c_{2}\tau_{1}\tau_{3}&-c_{1}\tau_{2}-c_{2}\tau_{1}\tau_{2}+c_{2}\tau_{3}&c_{0}+c_{1}\tau_{1}+c_{2}(\tau_{1}^{2}-\tau_{2})\end{array}\right)$
From Vieta’s formulas, we have $\tau_{1}=-a_{1},\tau_{2}=a_{2},\tau_{3}=-a_{3}.$ The second statement follows just by calculating $\det B$, so (i) follows.
(ii) From $BM_{P}=M_{P}B$, the expression of B in (i) follows. Or applying (i) with $a_{1}=a_{2}=0,a_{3}=-a.$
(iii) By (ii), $B\in Saut_{P}$ if and only if
$\det B=c_{0}^{3}+c_{1}^{3}+c_{2}^{2}-3c_{0}c_{1}c_{2}$
$=(c_{0}+c_{1}+c_{2})(c_{0}^{2}+c_{1}^{2}+c_{2}^{2}-c_{0}c_{1}-c_{1}c_{2}-c_{2}c_{1})$
$=\frac{1}{2}(c_{0}+c_{1}+c_{2})((c_{0}-c_{1})^{2}+(c_{1}-c_{2})^{2}+(c_{2}-c_{0})^{2})$
Therefore, by the second line above $\lambda_{1}\lambda_{2}\lambda_{3}=\pm 1$ holds only if $c_{0}+c_{1}+c_{2}=\pm 1$. Hence, by the third line above $(c_{0}-c_{1})^{2}+(c_{1}-c_{2})^{2}+(c_{2}-c_{0})^{2}=2$. This is possible only if two of the terms in the left hand side is 1 and the other is 0. Without loss of generality, assume that $c_{0}=c_{1}$ and $c_{2}=c_{0}\pm 1$. Then, we have $c_{0}+c_{1}+c_{2}=3c_{0}\pm 1$. Since $c_{0}+c_{1}+c_{2}=\pm 1$, we must have $c_{0}=0$, and hence, $(c_{0},c_{1},c_{2})=(0,0,\pm 1)$. Thus, $\lambda_{1}\lambda_{2}\lambda_{3}=\pm 1$ is equivalent to $(c_{0},c_{1},c_{2})\in\{(0,0,\pm 1),(0,\pm 1,0),(\pm 1,0,0)\}$, and hence, $(\lambda_{1},\lambda_{2},\lambda_{3})\in\{(\pm 1,\pm 1,\pm 1),(\pm\omega,\pm\omega,\pm\omega),(\pm\omega^{2},\pm\omega^{2},\pm\omega^{2})\}$. Thus, $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order 3. ∎
Example 7.
Let $P(z)=-z^{3}+1\in\mathbb{Z}[z]$ and $\omega=\frac{-1+\sqrt{3}}{2}$. Then, by Theorem 7, $B=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)$. Since we can easily verify that
$B^{2}=\left(\begin{array}[]{ccc}0&0&1\\ 1&0&0\\ 0&1&0\end{array}\right)\neq I_{3},~{}B^{3}=I_{3}.$
Therefore, $Saut_{P}/\{\pm I_{3}\}$ is a cyclic group of order $3$ and it is spanned by $B=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)$. Furthermore Let $X=(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $\varphi_{B}(X)=(y_{n};n\in\mathbb{Z})$. Then,
$\left(\begin{array}[]{c}y_{0}\\ y_{1}\\ y_{2}\end{array}\right)=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)\left(\begin{array}[]{c}x_{0}\\ x_{1}\\ x_{2}\end{array}\right)=\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ x_{0}\end{array}\right).$
Since $x_{3}=x_{0},~{}y_{3}=y_{0}$, we have
$\left(\begin{array}[]{c}y_{1}\\ y_{2}\\ y_{3}\end{array}\right)=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ x_{3}\end{array}\right)=\left(\begin{array}[]{c}x_{2}\\ x_{3}\\ x_{1}\end{array}\right)=\left(\begin{array}[]{c}x_{2}\\ x_{0}\\ x_{1}\end{array}\right)$
as expected.
7 Streams over Galois fields
In the last section, We investigate the streams over the Galois field $GF(q^{k})=GF(q)(\theta)$, where $\theta$ is a root of an irreducible polynomial $P(z)\in GF(q)[z]$ of degree $k$. The zeros $\Omega_{P}^{q}$ of $ST(GF(q^{k}))$ by the action of $P(Z)$ in the above sense is discussed. Note that any element in $\Omega_{P}^{q}$ has cycle $q^{k}$, any automorphism of $\Omega_{P}^{q}$ is an strong automorphism, and hence, is represented by a $k\times k$ matrix with entries in $GF(q)$.
Theorem 8.
Let $P(z)\in GF(q)[z]$ be an irreducible polynomial of degree $k\geq 1$. Let $\Omega_{P}^{q}$ be the zeros of the $P(z)$-action to $St(GF(q))$. Then, the group of the automorphisms of $\Omega_{P}^{q}$ commuting with the shift is a cyclic group of order $q^{k}-1$.
Proof.
Let $P(z)=a_{0}+a_{1}z+\cdots+a_{k}z^{k}$ with $a_{i}\in GF(q)$. We may assume $a_{0}=1$ since $P(z)$ is irreducible and $GF(q)$ is a field. Then, any element in $\Omega_{P}^{q}$ is cyclic with period $q^{k}-1$, since for any $(x_{n};n\in\mathbb{Z})\in\Omega_{P}^{q}$, $x_{n}\in GF(q)~{}(\forall n\in\mathbb{Z})$ is determined by $(x_{n-k},\cdots x_{n-1})\in GF(q)^{k}$. Hence, any automorphism of $\Omega_{P}^{q}$ commuting with the shift can be written as a $k\times k$-matrix with entries in GF(q).
Let $\theta_{1},\cdots,\theta_{k}$ be the roots of $P(z)=0$. Let $\eta$ be a primitive element of $GF(q^{k})=GF(q)(\theta_{1},\cdots,\theta_{k})$. That is, $\eta^{i}=1$ if and only if $i\equiv 0~{}(mod~{}q^{k}-1)$. Then, $\eta$ can be written as $\eta=c_{1}\theta_{1}+\cdots+c_{k}\theta_{1}^{k}$ with $c_{1},\cdots,c_{k}\in GF(q)$ with $c_{1}\cdots c_{k}\neq 0$ Let
$\lambda_{i}=c_{1}\theta_{i}+\cdots+c_{k}\theta_{i}^{k}~{}(i=1,\cdots,k).$
Let the matrix $B$ be defined by (5). Then, it is a matrix with entries in $GF(q)$ since the entries are symmetric functions of $\theta_{1},\cdots,\theta_{k}$. Also, it commutes with $M_{P}$. Therefore, it gives an automorphism of $\Omega_{P}^{q}$ commuting with the shift. Moreover, since $\lambda_{1}=\eta$ has the multiplicative order $q^{k}-1$, we have $B^{i}=I_{k}$ if and only if $i\equiv 0~{}(mod~{}q^{k}-1)$. This implies that the set of automorhisms of $\Omega_{P}^{q}$ commuting with the shift is a cyclic group of order $q^{k}-1$, since the cardinality of the set of $B$ as (5) which is nonsingular is $q^{k}-1$. ∎
Example 8.
Let us consider $P(z)=z^{2}+1\in GF(3)[z]$. Let ±$\pm\theta\in GF(9)$ be the roots of $z^{2}+1=0$. Then, ±$\pm\theta+1$ are the primitive elements of $GF(9)$. Then, the following $B$ generates the automorphism group of $\Omega_{P}^{3}$ commuting with the shift.
$B=\left(\begin{array}[]{cc}2&2\\ \theta&-\theta\end{array}\right)\left(\begin{array}[]{cc}\theta+1&0\\ 0&-\theta+1\end{array}\right)\left(\begin{array}[]{cc}2&2\\ \theta&-\theta\end{array}\right)^{-1}=\left(\begin{array}[]{cc}2&1\\ 2&2\end{array}\right)$
$012$
Figure 2: $\sigma$-orbits of $\Omega_{P}^{3}$
012012
Figure 3: $B$-orbit of $\Omega_{P}^{3}$
The $\sigma$-orbits of $\Omega_{P}^{3}$ and the $B$-orbit of $\Omega_{P}^{3}$ are as above.
Acknowledgements
The authors thank Prof. Shigeki Akiyama and Prof. Hajime Kaneko (Tsukuba Univ., Japan) for their useful suggestions. The simple proof of (iii) of Theorem 7 was taught by Prof. Shigeki Akiyama. The first author was supported by TianYuan Visiting Scholar Program of NSFC(Grant No:12426661).