### v5

# Dynamical systems defined by polynomials with algebraic properties

Abstract
Let $St(G)$ be the set of streams over an additive group $G$, that is, the set of two side sequences from $\prod_{-\infty}^{\infty}G$. We are mainly interested in the case that $G$ is the torus $\mathbb{T}=\mathbb{R}/\mathbb{Z}$. Given a primitive polynomial $P(z)\in\mathbb{Z}[z]$ and identifying $(x_{n};n\in\mathbb{Z})\in St(\mathbb{T})$ as the formal power series $\sum_{n\in\mathbb{Z}}x_{n}z^{n}$, we shall study algebraic and dynamical properties of the space of stream zeros $\Omega_{P}\subset St(\mathbb{T})$, which is the kernel of the action $P(z)$ on $St(\mathbb{T})$, where the action is the multiplication as formal power series over $z$, considering $X:=(x_{n};n\in\mathbb{Z})\in St(\mathbb{T})$ as the formal power series $\sum_{n\in\mathbb{Z}}x_{n}z^{n}$. Together with the shift $\sigma$ defined by $(\sigma X)_{n}=x_{n+1}~(\forall n\in\mathbb{Z})$, $\Omega_{P}$ becomes a compact dynamical system $(\Omega_{P},\sigma)$ whose symbolic representation is given by the integer part of the image of the action. We proved that there is a one-to-one correspondence between the factorization of $P(z)$ and the semi-direct sum decomposition of the additive group $\Omega_{P}$. Furthermore, the group of strong automorphisms of $\Omega_{P}$ is investigated. Specially in the case of $\deg P=2$, it is determined by the discriminant of $P(z)$ and can be represented by the continued fraction expansion of its roots. Finally, we study the streams over the Galois field $GF(q)$. In this case, the group of strong automorphisms of $\Omega_{P}^{q}$ is a cyclic group of order $q^{k}-1$, where $k$ is the degree of $P(z)$.
1 Introduction
Let $X=(x_{n}\in G;n\in\mathbb{Z})$ be a two-side sequence of elements in an additive group $G$, which is called a two side stream (or simply, stream) over $G$. The set of two side streams over $G$ is denoted by $St(G)$. This $X$ is written as $(\cdots,x_{-2},x_{-1};x_{0},x_{1},\cdots)$, here the semicolon “;” separates the coordinates $x_{n}~(n<0)$ from $x_{n}~(n\geq 0)$. It is considered also as a two-side formal power series in a variable $z$, namely,
$X=\cdots+x_{-2}z^{-2}+x_{-1}z^{-1}+x_{0}+x_{1}z+x_{2}z^{2}+\cdots.$
Throughout this paper, we use the variable $z$ in this sense.
$St(G)$ is considered as an additive group by component-wise addition. Moreover, if $G$ is a topological group, then $St(G)$ is considered as a topological group with the product topology coming from $G$. Let 0 be the additive unit of $G$, then the additive unit of $St(G)$ is $(0^{\infty};0^{\infty})$, where $0^{\infty}$ before “;” implies $\cdots,0,0$ while that after “;” implies $0,0,\cdots$.
Let $G$ be a topological commutative ring having the multiplication unit 1. The product of $(x_{n};n\in\mathbb{Z})$ and $(y_{n};n\in\mathbb{Z})\in St(G)$ as the formal power series is defined if it exists, and called convolution product as well, thta is
$(x_{n};n\in\mathbb{Z})\times(y_{n};n\in\mathbb{Z})=\left(\sum_{i=-\infty}^{\infty}x_{i}y_{n-i};~n\in\mathbb{Z}\right).$ (1)
The unit of $St(G)$ with respect to the convolution product is $I=(0^{\infty};1,0^{\infty})$. We call $(x_{n};n\in\mathbb{Z})\in St(G)$ absolutely summable if the series $\sum_{n\in\mathbb{Z}}x_{n}$ converges to the same value independent of the order of the sum.
One-sided sequences of $G$ have been studied for a long time in the field of computer science as stream calculus (see [5], for example), where the convolution product always exists since the above sum becomes a finite sum. In this paper, we study two side sequences and call them streams.
For $(x_{n};n\in\mathbb{Z})\in St(G)$, even if $(y_{n};~n\in\mathbb{Z})\in St(G)$ such that
$(x_{n};n\in\mathbb{Z})\times(y_{n};~n\in\mathbb{Z})=I$
exists, it may not be unique. For example,
$(0^{\infty};1,-1,0^{\infty})\times(0^{\infty};1^{\infty})=(0^{\infty};1,-1,0^{\infty})\times((-1)^{\infty};0^{\infty})=I.$
An element $(a_{n};n\in\mathbb{Z})$ of $St(\mathbb{C})$ such that $a_{n}=0$ except for finitely many $n$ can be identified with a polynomial $P(z)=\sum_{n\in S}a_{n}z^{n}$ admitting negative powers, where $S\subset\mathbb{Z}$ is the finite set of $n$ such that $a_{n}\neq 0$.
Definition 1.
We call a polynomial $P(z)=\sum_{n\in S}a_{n}z^{n}\in\mathbb{Z}[z]$ primitive if $S\subset\{0,1,\cdots,k\}$ with $k\geq 1$, $a_{0}\neq 0,a_{k}\neq 0$ and the greatest common divisor (GCD) of the coefficients $a_{0},a_{1},\cdots,a_{k}$ is $\pm 1$. We call $P(z)$ hyperbolic if $P(z)=0$ does not hold for any $z\in\mathbb{C}$ with $|z|=1$.
Let $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ be the torus, which is a topological additive group with $\mathbb{Z}$-action. Hence, $St(\mathbb{T})$ is considered as a topological additive group with $\mathbb{Z}$-action, where the $\mathbb{Z}$-action is coodinatewise. Let $P(z)\in\mathbb{Z}[z]$ be a primitive polynomial as in Definition 1. Then, $P(z)$ acts on $St(\mathbb{T})$ as the multiplication of formal power series or the convolution product (1). That is,
$(P(z),(x_{n};n\in\mathbb{Z}))\mapsto(a_{n};n\in S)\times(x_{n};n\in\mathbb{Z})\\ =(a_{k}x_{n-k}+\cdots+a_{1}x_{n-1}+a_{0}x_{n};n\in\mathbb{Z})\in St(\mathbb{T}).$
We study the kernel of this action, denoted by $\Omega_{P}$ and is called the stream zeros of $P(z)$, that is
$\Omega_{P}:=\left\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});P(z)\times(x_{n};n\in\mathbb{Z})=(0^{\infty};0^{\infty})\right\}.$
We show that $\Omega_{P}$ has a lot of similarities with the algebraic roots of $P(z)=0$.
The paper is organized as follows. In Section 2 we show some basic properties of the convolution product which is useful for further sections. In Section 3 we discuss $\Omega_{P}$ from dynamical system viewpoint. In Section 4 we introduce the notion of lift of a steam zeros on $\mathbb{T}$, that is, a stream zeros on $\mathbb{R}$ whose projection becomes it and prove Theorem 3 which gives a way to get a lift if Condition (#) is satisfied. Moreover, we prove that any primitive $P(z)$ satisfies Condition (#), which is of independent interest. When the lifting matrix of size $l=1$, this was answered by M. Newman [3] and X. Zhan [11]. In Section 5 we use the tools from resultant of two polynomials, generalized Vandermonde’s determinant, and Hadamard product to establish one-to-one correspondence between the factorization of $P(z)$ and the semi-direct sum decomposition of the additive group $\Omega_{P}$. In Section 6, we turn to investigate the algebraic structure of $\Omega_{P}$, we prove that the group of strong automorphisms of $\Omega_{P}$ commuting with the shift can be represented by the matrices from $Saut_{P}$ (Definition 4). In Section 7, we discuss the case $\deg P(z)=2$ and show that $Saut_{P}$ is determined by the discriminant of $P(z)$, while when the degree is larger than 2, we obtain only partial results, which also indicates that it has close relationship with the theory of Diophantine equations, known as Pell’s equation shown in the former case. We provide some examples to illustrate all theses theorem. In the last section, we focus on the zeros of $P(z)$ action on $St(GF(q))$, where $GF(q)$ is the Galois field of size $q$. We prove that if $P(z)$ is irreducible, then $Saut_{P}$ is a cyclic group of order $q^{k}-1$.
2 Convolution product and its inverse
In this section, we present some basic properties of the convolution product of $St(\mathbb{C})$. As one can see from the previous discussion, the convolution inverse may not be unique, we thus wish to find some conditions that can guarantee the uniqueness of the convolution inverse. The following theorem is known ([8]), here we give another elementary proof.
Theorem 1.
Let $U=(u_{n};n\in\mathbb{Z}),~V=(v_{n};n\in\mathbb{Z}),~W=(w_{n};n\in\mathbb{Z})$ be arbitrary elements in $St(\mathbb{C})$.
(i) We have $U\times V=V\times U$ if one of $U\times V$ or $V\times U$ exists. If among $U,V,W$, two of them are absolutely summable and the remaining one is bounded, then the convolution products below exist and it holds that
$(U\times V)\times W=U\times(V\times W).$
(ii) If $U$ is absolutely summable, then an absolutely summable $V$ satisfying $U\times V=I$ or $V\times U=I$ is unique if exists. In this case, we have $U\times V=V\times U=I$ and We call $V$ as this convolution inverse of $U$ and is denoted by $U^{-1}$.
(iii) If both of $U,V\in St(\mathbb{C})$ are absolutely summable, then so is $U\times V$. Furthermore, if both of $U^{-1},V^{-1}$ exist, then $(U\times V)^{-1}$ exists and $(U\times V)^{-1}=V^{-1}\times U^{-1}$ holds. Moreover, $(U^{-1})^{-1}=U$ holds.
(iv) Any primitive polynomial $P(z)=(a_{n};n\in\mathbb{Z})$ has a unique absolutely summable inverse $P(z)^{-1}$ in $St(\mathbb{R})$ if it is hyperbolic.
Proof.
(i) The firsr statement is clear. For the second statement, let $U,V,W$ satisfy the required condition. Then for any $n\in\mathbb{Z}$, we prove that
$\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}|u_{i}v_{j}w_{k}|<\infty.$
By symmetry, we may assume that $U,V$ are absolutely summable and $W$ is bounded. Then, there exists a constant $L>0$ such that
$\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}|u_{i}v_{j}w_{k}|<L\sum_{i,j\in\mathbb{Z}}|u_{i}v_{j}|=L\sum_{i\in\mathbb{Z}}|u_{i}|\sum_{j\in\mathbb{Z}}|v_{j}|<\infty.$
Thus, $\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}u_{i}v_{j}w_{k}$ converges absolutely. Hence, we have
$\sum_{k\in\mathbb{Z}}\left(\sum_{i,j\in\mathbb{Z}\atop{i+j=n-k}}u_{i}v_{j}\right)w_{k}=\sum_{i\in\mathbb{Z}}u_{i}\left(\sum_{j,k\in\mathbb{Z}\atop{j+k=n-i}}v_{j}w_{k}\right)=\sum_{i,j,k\in\mathbb{Z}\atop{i+j+k=n}}u_{i}v_{j}w_{k}$
and $(U\times V)\times W=U\times(V\times W)$.
(ii) We prove the uniqueness since the other statements follow from the first statement in (i). If $U\times V=U\times V^{\prime}=I$, then by (i), we have
$V^{\prime}=(UV)V^{\prime}=(VU)V^{\prime}=V(UV^{\prime})=V.$
(iii) is clear.
(iv) Let $P(z)=\sum_{n\in S}a_{n}z^{n}$ with a finite nonempty set $S=\{n\in\mathbb{Z};a_{n}\neq 0\}$. Consider the factorization of $P(z)$ and applying (ii), it is sufficient to prove that $P(z)$ being of the form, either constant $C\neq 0$, $z,z^{-1}$ or $z-\omega$ with $0<|\omega|\neq 1$, has the absolutely summable inverse.
Case 1: $P(z)=C$. Since $C=(0^{\infty};C,0^{\infty})$, $C^{-1}=(0^{\infty};c^{-1},0^{\infty})$ becomes the absolutely summable inverse.
Case 2: $P(z)=z~\mbox{or}~z^{-1}$. One easily verifies that $z$ and $z^{-1}$ are absolutely summable and inverses of each other. That is, $(0^{\infty}1;0^{\infty})=(0^{\infty};0,1,0^{\infty})^{-1}$.
Case 3: P(z)=zwP(z)=z-w with $|\omega|>1.$ Since
$(z-\omega)^{-1}=(-1)\omega^{-1}(1-\omega^{-1}z)^{-1}=-\omega^{-1}-\omega^{-2}z-\omega^{-3}z^{2}-\cdots,$
so $(0^{\infty};-\omega^{-1},-\omega^{-2},-\omega^{-3},\cdots)$ becomes the absolutely summable inverse of $z-\omega$.
Case 4: P(z)=zwP(z)=z-w with $|\omega|<1$. Since
$(z-\omega)^{-1}=z^{-1}(1-\omega z^{-1})^{-1}=z^{-1}+\omega z^{-2}+\omega^{2}z^{-3}+\cdots,$
then $(\cdots,\omega^{2},\omega,1;0^{\infty})$ becomes the absolutely summable inverse of $z-\omega$.
If $\omega\notin\mathbb{R}$ in Case 2 (or Case 3), then the conjugate $\overline{\omega}$ is also in Case 2 (or Case 3, respectively) and $(z-\omega)^{-1}(z-\overline{\omega})^{-1}\in St(\mathbb{R})$. Thus, $P(z)$ has the absolutely summable inverse $P(z)^{-1}\in St(\mathbb{R})$ as the product of the above terms. ∎
3 Stream zeros and dynamical systems
Let $P(z)$ be a primitive polynomial over $\mathbb{Z}$. Then clearly $\Omega_{P}$ is a compact $\mathbb{Z}$-subgroup of $St(\mathbb{T})$. Moreover, the pair $(\Omega_{P},\sigma)$, where $\sigma$ is the shift, becomes a compact dynamical system, which is called the dynamical system introduced by $P(z)$. On the other hand, let
$\{\Omega_{P}\}=\{(\{x_{n}\};n\in\mathbb{Z})\in St([0,1));~(x_{n};n\in\mathbb{Z})\in\Omega_{P}\},$
where $\{x_{n}\}$ is the representative of $x_{n}\in\mathbb{T}$ in $[0,1)$. Then, $P(z)\times X\in St(K_{P})$ holds for any $X\in\{\Omega_{P}\}$, where $K_{P}$ is a finite subset of integers with
$K_{P}=\{k_{*}+1,k_{*}+2,\cdots,k^{*}-1\},~\mbox{where}$
[…]
Let
$\Xi_{P}=\{P(z)\times\{\tilde{x}_{n}\};~\{\tilde{x}_{n}\}\in\{\Omega_{P}\}\}.$ (3)
Then $\Xi_{P}$ is a compact, shift-invariant subset of $K_{P}^{\mathbb{Z}}$. Together with the shift $\sigma$, we call $(\Xi_{P},\sigma)$ the symbolic representation of $(\Omega_{P},\sigma)$
Theorem 2.
Let a primitive polynomial $P(z)=\sum_{n\in S}a_{n}z^{n}$ be hyperbolic. Then, the dynamical systems $(\Omega_{P},\sigma)$ and $(\Xi_{P},\sigma)$ are conjugate. Hence, we have the following commutative diagram. The conjugate map $\times P(z)$ is almost continuous, that is, continuous except for a countable set.
Moreover, a necessary and sufficient condition for $\xi=(\xi(n);~n\in\mathbb{Z})\in St(\mathbb{Z})$ to be $\xi\in\Xi_{P}$ is that $P(z)^{-1}\times\xi\in St([0,1))$, which automatically implies that $P(z)^{-1}\times\xi\in\Omega_{P}$.
$\{\Omega_{P}\}$$\{\Omega_{P}\}$$\Xi_{P}$$\Xi_{P}$$\sigma$$\sigma$$\times P(z)^{-1}$$\times P(z)^{-1}$$\times P(z)$$\times P(z)$
Proof.
The first half of the statement is clear. We prove the second half. If $\xi\in\Xi_{P}$, then there exists $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ such that $\xi=P(z)\times(\{x_{n}\};n\in\mathbb{Z})$. Therefore, we have
$P(z)^{-1}\times\xi=P(z)^{-1}\times P(z)\times(\{x_{n}\};n\in\mathbb{Z})=(\{x_{n}\};n\in\mathbb{Z})\in[0,1)^{\mathbb{Z}}.$
Conversely, assume that $y:=P(z)^{-1}\times\xi\in[0,1)^{\mathbb{Z}}$ and $x=\pi(y)\in St(\mathbb{T})$, $\pi:\mathbb{R}\to\mathbb{T}$ being the projection. Since $P(z)\times y=\xi\in K_{P}^{\mathbb{Z}}\subset\mathbb{Z}^{\mathbb{Z}}$, we have $P(z)\times x=(0^{\infty};0^{\infty})\in St(\mathbb{T})$, and hence, $x\in\Omega_{P}$. Hence, $\xi=P(z)\times y=P(z)\times\{x\}$ for some $x\in\Omega_{P}$. Thus, $\xi\in\Xi_{P}$. ∎
Remark 1.
To recover a dynamical system on $St([0,1))$ from its symbolic representation is known as Kaneko’s formula [8]. It is nothing but to find an absolutely summable $P(z)^{-1}$ as shown in Theorem 2.
Example 1.
Let $P(z)=z^{2}-3z+1$. Then, we have
$\Omega_{P}=\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});~x_{n+2}=3x_{n+1}-x_{n}~\mbox{(mod 1) for any}~n\in\mathbb{Z}\}.$
We rewrite the above as
$\Omega_{P}=\left\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});~\left(\begin{array}[]{c}x_{n+1}\\ x_{n+2}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ -1&3\end{array}\right)\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\end{array}\right)\mbox{(mod 1) for any }n\in\mathbb{Z}\right\}.$
Observe that
$P(z)=z^{2}-3z+1=(z-\omega_{+})(z-\omega_{-}),\quad\mbox{here}~\omega_{\pm}:=\frac{3\pm\sqrt{5}}{2}.$
Since $|\omega_{+}|=|\frac{3+\sqrt{5}}{2}|>1$,
$(z-\omega_{+})^{-1}=(-1)\omega_{+}^{-1}(1-\omega_{+}^{-1}z)^{-1}=-\omega_{+}^{-1}-\omega_{+}^{-2}z-\omega_{+}^{-3}z^{2}-\cdots,$
Similarly since $|\omega_{-}|=|\frac{3-\sqrt{5}}{2}|<1$,
$(z-\omega_{-})^{-1}=z^{-1}(1-\omega_{-}z^{-1})^{-1}=z^{-1}+\omega_{-}z^{-2}+\omega_{-}^{2}z^{-3}+\cdots.$
Since
$\omega_{+}^{-1}\sum_{n=0}^{\infty}(\omega_{+}^{-1}\omega_{-})^{n}=\omega_{+}^{-1}\frac{1}{1-\omega_{+}^{-1}\omega_{-}}=\frac{1}{\omega_{+}-\omega_{-}}=\frac{1}{\sqrt{5}},$
we have the following Kaneko’s formula:
$P(z)^{-1}=(z-\omega_{+})^{-1}\times(z-\omega_{-})^{-1}$
$=(0^{\infty};-\omega_{+}^{-1},-\omega_{+}^{-2},-\omega_{+}^{-3},\cdots)\times(\cdots,\omega_{-}^{3},\omega_{-}^{2},\omega_{-},1;0^{\infty})$
$=-\frac{1}{\sqrt{5}}(\cdots,\omega_{-}^{3},\omega_{-}^{2},\omega_{-},1;\omega_{+}^{-1},\omega_{+}^{-2},\omega_{+}^{-3},\cdots).$
It is interesting that the above formula can also be obtained by the so-called partial fractional decomposition.
$P(z)^{-1}=(z-\omega_{+})^{-1}\times(z-\omega_{-})^{-1}=\frac{1}{\omega_{+}-\omega_{-}}\{\frac{1}{z-\omega_{-}}-\frac{1}{z-\omega_{-}}\}$
$=-\frac{1}{\sqrt{5}}\left\{\frac{1}{\omega_{+}}\frac{1}{1-z\omega_{+}^{-1}}+\frac{1}{z}\frac{1}{1-z^{-1}\omega_{-}}\right\}$
$=-\frac{1}{\sqrt{5}}\left\{\frac{1}{\omega_{+}}\sum_{n=0}^{\infty}(\omega_{+}^{-1}z)^{n}+\frac{1}{z}\sum_{n=0}^{\infty}(\omega_{-}z^{-1})^{n}\right\}$
$=-\frac{1}{\sqrt{5}}(\cdots,\omega_{-}^{3},\omega_{-}^{2},\omega_{-},1;\omega_{+}^{-1},\omega_{+}^{-2},\omega_{+}^{-3},\cdots).$
$x_{n}$$x_{n+1}$1101-12-2
Figure 1: Symbolic representation of $St([0,1))$: $\xi_{n+2}\in\{-2,-1,0,1\}.$
Identifying $\mathbb{T}$ with $[0,1)$, $\Omega_{P}$ can be considered as the set of trajectories of the dynamical system on $0,1)\times[0,1)$ with the transformation $(x,y)\mapsto(y,3y-x)~(\rm{mod}~1)$. That is, $x_{n+2}=\{3x_{n+1}-x_{n}\}~(\forall n\in\mathbb{Z})$ (remind that $\{~\}$ is the fractional part). The symbolic representation of $(x_{n}\in[0,1);n\in\mathbb{Z})$ is the sequence $(\xi_{n}\in K_{P};n\in\mathbb{Z})$ with $K_{P}=\{-2,-1,0,1\}$ and $\xi_{n+2}=x_{n}-3x_{n+1}+x_{n+2}\in K_{P}~(\forall n\in\mathbb{Z})$ (Figure 1).
The topological entropy of these dynamical systems coincides with
$\lim_{n\to\infty}\frac{1}{n}\log\#\Xi_{P}|_{\{1,2,\cdots,n\}}.$
On the other hand, it is well known that it coincides with the sum of logarithms of the eigenvalues greater than 1 of the companion matrix (see [9], for example). In our case, the companion matrix $\left(\begin{array}[]{cc}0&1\\ -1&3\end{array}\right)$ has eigenvalues $\frac{3\pm\sqrt{5}}{2}$, and hence, the topological entropy is $\log\frac{3+\sqrt{5}}{2}$.
4 Lifts of stream zeros
Let $P(z)=a_{k}z^{k}+a_{k-1}z^{k-1}+\cdots+a_{0}\in\mathbb{Z}[z]$ be a primitive polynomial. Let $N\in\mathbb{Z}_{>0}$. Define
$\tilde{\Omega}_{P}:=\{(\tilde{x}_{n}\in\mathbb{R};~n\in\mathbb{Z});~a_{k}\tilde{x}_{n-k}+\cdots+a_{k-1}\tilde{x}_{n-k+1}+a_{0}\tilde{x}_{n}=0~(\forall n\in\mathbb{Z})\}$
and its restriction to [N,N][-N,N]
$\tilde{\Omega}_{P}|_{N}:=\{(\tilde{x}_{n}\in\mathbb{R};~n\in[-N,N]);~(\tilde{x}_{n}\in\mathbb{R};~n\in\mathbb{Z})\in\tilde{\Omega}_{P}\}.$
For $l=1,2,\cdots$, define a $l\times(k+l)$-matrix $M_{P}^{(l)}$
$M_{P}^{(l)}=\left(\begin{array}[]{ccccccccc}a_{k}&a_{k-1}&\cdot&a_{0}&0&\cdot&\cdot&\cdot&0\\ 0&a_{k}&a_{k-1}&\cdot&a_{0}&0&\cdot&\cdot&0\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&0\\ \cdot&\cdot&\cdot&\cdot&0&a_{k}&a_{k-1}&\cdot&a_{0}\end{array}\right).$ (4)
Definition 2.
We say that $(x_{n};n\in\mathbb{Z})\in St(\mathbb{T})$ has a lift in $\tilde{\Omega}_{P}$ if for any $N\in\mathbb{Z}_{>0}$, there exists $(\tilde{x}_{n}\in\mathbb{R};~-N\leq n\leq N)\in\tilde{\Omega}_{P}|_{N}$ such that $\pi(\tilde{x}_{n})=x_{n}$ for any $n\in[-N,N]$, where $\pi:\mathbb{R}\to\mathbb{T}$ is the the projection. We call $(\tilde{x}_{n}\in\mathbb{R};~-N\leq n<N)$ as above a local lift of $(x_{n};n\in\mathbb{Z})$ in $\tilde{\Omega}_{P}$. We say that $\Omega_{P}$ has a lift if any of $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ has a lift in $\tilde{\Omega}_{P}$.
The following Lemma due to S. Akiyama [12] is essential to prove Theorem 3.
Lemma 1.
[12] In the hyperplane $H$ of $\mathbb{R}^{k+l}$ spanned by the row vectors of $M_{P}^{(l)}$, any integer point in $H$ can be written as a linear combination of the row vectors of $M_{P}^{(l)}$ with integer cooeffficients. Hence, the parallelpiped spanned by the row vectors of $M_{P}^{(l)}$ contains integer points only at the vertices.
Proof.
We first prove that if $\eta=(x_{1},x_{2},\cdots,x_{l})$ is a row vector of degree $k$ with real entries such that $\eta M_{P}^{(l)}$ is an integer vector, then $\eta$ is an integer vector.
We may assume that $\eta$ does not contain irrational elements. Suppose to the contrary that it contains irrational elements and $x_{i}$ is one of them with the minimum $i$. Then the $i$-the element of $\eta M_{P}^{(l)}$ is irrational, contradicting the assumption. Hence, we may assume that $\eta$ is a rational vector. Furthermore, it is sufficient to prove that for any prime number $p$, $\nu_{p}(x_{i})\geq 0~(i=1,\cdots,l)$, where $\nu_{p}$ is the discrete valuation at $p$.
Suppose to the contrary that $\nu_{p}(x_{i})<0$ for some $p$ and $i$. Let the minimum $\nu_{p}(x_{i})$ as this be $m_{0}$. Let $i_{0}$ be such that
$\nu_{p}(x_{i_{0}})=m_{0}<\nu_{p}(x_{i})~(1\leq\forall i<i_{0})~\mbox{and}~\nu_{p}(x_{i_{0}})\leq\nu_{p}(x_{i})~(i_{0}<\forall i\leq l).$
Since the greatest common devisor of $a_{k},\cdots,a_{1},a_{0}$ is 1, there exists $a_{j}$ such that $\nu_{p}(a_{j})=0$. Let $j_{0}$ be such that
$0=\nu_{p}(a_{j_{0}})<\nu_{p}(a_{j})~(0\leq\forall j<j_{0}).$
Then, let the $(k+i_{0}-j_{0})$-th element of $\eta M_{P}^{(l)}$ be $\xi$, then since
$\xi=x_{1}a_{j_{0}-i_{0}+1}+\cdots+x_{i_{0}}a_{j_{0}}+\cdots+x_{l}a_{j_{0}-i_{0}+l}~(\mbox{we put}~a_{j}=0~\mbox{if}~j\notin\{0,1,\cdots,k\})$
and $\nu_{p}(x_{i}a_{j_{0}-i_{0}-i})>\nu(x_{i_{0}}a_{j_{0}})~(\forall i\neq i_{0})$, we have
$\nu_{p}(\xi)=\nu_{p}(x_{i_{0}})+\nu_{p}(a_{j_{0}})<0.$
Thus, $\xi$ is not an integer.
This implies that any integer point $(y_{0},y_{1},\cdots,y_{k+l})$ in the space spanned by the row vectors of $M_{P}^{(l)}$ is an integer combination of the row vectors of $M_{P}^{(l)}$. ∎
Definition 3.
We say that a primitive polynomial $P(z)=a_{k}z^{k}+a_{k-1}z^{k-1}+\cdots+a_{0}\in\mathbb{Z}[z]$ satisfies Condition (#) if for any $l=1,2,\cdots$, there exists a $k\times(k+l)$-matrix $R$ whose entries are integers such that $\det\left(\begin{array}[]{c}M_{P}^{(l)}\\ R\end{array}\right)=\pm 1$. Hence, $\left(\begin{array}[]{c}M_{P}^{(l)}\\ R\end{array}\right)$ has the inverse whose entries are integers. Let it be $(\tilde{L},L)$, where $\tilde{L},L$ are $(k+l)\times l$-matrix and $(k+l)\times k$-matrix. Let $\Lambda=LR$. It is a square matrix of size $k+l$ called a lifting matrix of $P(z)$ of size $l$.
Theorem 3.
A primitive, polynomial $P(z)$ satisfies Condition (#) always, and hence, $\Omega_{P}$ has a lift. In fact, for any large $N\in\mathbb{Z}_{>0}$ and l=2Nk+1l=2N-k+1, let $\Lambda$ be a lifting matrix of size $l$ as in Definition $3$. Then, for any $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$, $\Lambda\left(\begin{array}[]{c}x_{-N}\\ x_{-N+1}\\ \vdots\\ x_{N}\end{array}\right)$ is a local lift of $(x_{n};-N\leq n\leq N)$ in $\tilde{\Omega}_{P}$.
Corollary 1.
(i) For any large $N$, let $\pi:\mathbb{R}^{\{N,-N+1,\cdots,N\}}\to\mathbb{T}^{\{N,-N+1,\cdots,N\}}$ be the coordinatewise projection, then it gives a surjective homomorphism from $\tilde{\Omega}_{P}|_{N}$ to $\Omega_{P}|_{N}$ as the topological additive groups with $\mathbb{Z}$-action.
(ii) For any $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and an integer $m$, there exists $(y_{n};n\in\mathbb{Z})\in\Omega_{P}$ such that
$m(y_{n};n\in\mathbb{Z}):=(my_{n};n\in\mathbb{Z})=(x_{n};n\in\mathbb{Z}).$
Proof.
We prove Corollary 1 first using Theorem 3. (i) is clear except for “surjective”, which also holds since $\Omega_{P}$ has a lift. To prove (ii), let $X:=(x_{n}\in\mathbb{T};~n\in\mathbb{Z})\in\Omega_{P}$. By Theorem 3, $X$ has a local lift in $\tilde{\Omega}_{P}|_{N}$ for any large $N\in\mathbb{Z}$. That is, there exists $(\tilde{x}_{n}^{(N)}\in\mathbb{R};-N\leq n\leq N)\in\tilde{\Omega}_{P}|_{N}$ such that $\pi(\tilde{x}_{n}^{(N)})=x_{n}~(-N\leq\forall n\leq N)$. Then, $(\tilde{x}_{n}^{(N)}/m\in\mathbb{R};-N\leq n\leq N)\in\tilde{\Omega}_{P}|_{N}$. Let $y_{n}^{(N)}=\pi(\tilde{x}_{n}^{(N)}/m)\in\mathbb{T}~(\forall n\in[-N,N])$. Then, $y_{n}^{(N)}\in\Omega_{P}|_{N}$. Since $St(\mathbb{T})$ is compact, there exisis a limiting point, say $Y:=(y_{n}\in\mathbb{T};n\in\mathbb{Z})$ as $N\to\infty$ of $y_{n}^{(N)}\in\Omega_{P}|_{N}$. Then, it is clear that $Y\in\Omega_{P}$ and $mY=X$.
Proof of Theorem 3 : Let $l=1,2,\cdots$ be arbitrary. To prove the Condition (#), it is sufficient to prove that there exists a $(l+1)\times(k+1+l)$-matrix $R$ such that the set of row vectors of $\left(\begin{array}[]{c}M_{P}^{(l)}\\ R\end{array}\right)$ is linear independent and the parallelepiped spanned by them contains the integer points only at vertices, since then the volume of the parallelepiped is 1 and hence, $\det\left(\begin{array}[]{c}M_{P}^{(l)}\\ R\end{array}\right)=\pm 1$.
We construct $R$ inductively starting from the 1st row. Since by Lemma 1, $M_{P}^{(l)}$ has the property that the raws are linearly independent and the parallelepiped spanned by them, say $\Gamma_{0}$, contains the integer points only at vertices. Then, we add any integer vector, say $\gamma$, independent of the raw vectors of $M_{P}^{(l)}$. Let $\Gamma$ be the parallelpiped spanned by $\Gamma_{0}$ and $\gamma$. Since both $\Gamma_{0}$ and $\gamma+\Gamma_{0}$ contains the integer points only at vertices, all the integer points in $\Gamma$ which are not the vertices are properly in between $\Gamma_{0}$ and $\gamma+\Gamma_{0}$. Let $\gamma^{\prime}$ be any one of them and $\Gamma^{\prime}$ be the parallelpiped spanned by $\Gamma_{0}$ and $\gamma^{\prime}$. Then the volumn of $\Gamma^{\prime}$ and the number of integer points which are not the vertices is strictly smaller than those of $\Gamma$. By this process, we finally find an integer vector $\gamma_{1}$ independent of the raw vectors of $M_{P}^{(l)}$ such that the parallelpiped $\Gamma_{1}$ spanned by $\Gamma_{0}$ and $\gamma_{1}$ has the property that all the integer points in it are only those at the vertices.
Repeating this process, we find a $(k+l)$-dimensioal parallelpipped $\Gamma_{k}$ in $\mathbb{R}^{k+l}$ which is spanned by $\Gamma_{0}$ and $\gamma_{1},\gamma_{2},\cdots,\gamma_{k}$ such that there are no integer points in $\Sigma_{k}$ other than the vertices. Thus, the volumn of $\Sigma_{k}$ is $1$ and $\det\left(\begin{array}[]{c}M_{P}^{(l)}\\ R\end{array}\right)=\pm 1$, where $R=\left(\begin{array}[]{c}\xi_{1}\\ \vdots\\ \xi_{k}\end{array}\right)$.
Let $N\in\mathbb{Z}_{>0}$ be large enough. Let l=2Nk+1l=2N-k+1. Now we prove that for any $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$, $\Lambda\left(\begin{array}[]{c}x_{-N}\\ x_{-N+1}\\ \vdots\\ x_{N}\end{array}\right)$ is a local lift of $(x_{n};-N\leq n\leq N)$ in $\tilde{\Omega}_{P}$. Assume without loss of generality that $x_{n}\in[0,1)~(\forall n\in\mathbb{Z})$. Let
$\xi_{n}=a_{k}x_{n-k}+a_{k-1}x_{n-k+1}+\cdots+a_{0}x_{n}\in\mathbb{Z}~~(\forall n\in\mathbb{Z}).$
Let $A=\left(\begin{array}[]{c}M_{P}^{(l)}\\ R\end{array}\right)$ be as in Definition 3. Then, we have
$A\left(\begin{array}[]{c}x_{-N}\\ x_{-N+1}\\ \vdots\\ x_{-N+l-1}\\ x_{-N+l}\\ \vdots\\ x_{-N+l+k-1}\end{array}\right)=\left(\begin{array}[]{c}\xi_{-N+k}\\ \xi_{-N+k+1}\\ \vdots\\ \xi_{-N+k+l-1}\\ \theta_{1}\\ \vdots\\ \theta_{k}\end{array}\right),$ (5)
where $\theta_{1},\cdots,\theta_{k}\in\mathbb{R}$ are not necessarily be specified. Let
$\left(\begin{array}[]{c}\eta_{-N}\\ \eta_{-N+1}\\ \vdots\\ \eta_{-N+l-1}\\ \eta_{-N+l}\\ \vdots\\ \eta_{-N+l+k-1}\end{array}\right):=A^{-1}\left(\begin{array}[]{c}\xi_{-N+k}\\ \xi_{-N+k+1}\\ \vdots\\ \xi_{-N+k+l-1}\\ 0\\ \vdots\\ 0\end{array}\right).$
Note that $\eta_{-N},\eta_{-N+1},\cdots,\eta_{-N+l+k-1}$ are integers and N+l+k1=N-N+l+k-1=N. Then, it follows that
$A\left(\begin{array}[]{c}x_{-N}-\eta_{-N}\\ x_{-N+1}-\eta_{-N+1}\\ \vdots\\ x_{-N+l-1}-\eta_{-N+l-1}\\ x_{-N+l}-\eta_{-N+l}\\ \vdots\\ x_{N}-\eta_{N}\end{array}\right)=\left(\begin{array}[]{c}0\\ 0\\ \vdots\\ 0\\ \theta_{1}\\ \vdots\\ \theta_{k}\end{array}\right),$
and hence,
$M_{P}^{(l)}\left(\begin{array}[]{c}x_{-N}-\eta_{-N}\\ x_{-N+1}-\eta_{-N+1}\\ \vdots\\ x_{N}-\eta_{N}\end{array}\right)=\left(\begin{array}[]{c}0\\ 0\\ \vdots\\ 0\end{array}\right).$
This implies that $(x_{n}-\eta_{n};-N\leq n\leq N)$ is a local lift of $(x_{n};-N\leq n\leq N)$ in $\tilde{\Omega}_{P}$.
Note that
$\left(\begin{array}[]{c}x_{-N}-\eta_{-N}\\ x_{-N+1}-\eta_{-N+1}\\ \vdots\\ x_{-N+l-1}-\eta_{-N+l-1}\\ x_{-N+l}-\eta_{-N+l}\\ \vdots\\ x_{N}-\eta_{N}\end{array}\right)=A^{-1}\left(\begin{array}[]{c}0\\ 0\\ \vdots\\ 0\\ \theta_{1}\\ \vdots\\ \theta_{k}\end{array}\right)=L\left(\begin{array}[]{c}\theta_{1}\\ \vdots\\ \theta_{k}\end{array}\right)$
and by (5),
$\left(\begin{array}[]{c}\theta_{1}\\ \theta_{2}\\ \vdots\\ \theta_{k}\end{array}\right)=R\left(\begin{array}[]{c}x_{-N}\\ x_{-N+1}\\ \vdots\\ x_{N}\end{array}\right).$
Therefore,
$\left(\begin{array}[]{c}x_{-N}-\eta_{-N}\\ x_{-N+1}-\eta_{-N+1}\\ \vdots\\ x_{N}-\eta_{N}\end{array}\right)=LR\left(\begin{array}[]{c}x_{-N}\\ x_{-N+1}\\ \vdots\\ x_{N}\end{array}\right)=\Lambda\left(\begin{array}[]{c}x_{-N}\\ x_{-N+1}\\ \vdots\\ x_{N}\end{array}\right).$
Remark 2.
In [12], the existence of the lift is proved directly. Here, we made a detour via the Condition (#)(\#), since it gives a concrete way to obtain it.
Lemma 2.
It holds that
$\det\left(\begin{array}[]{ccccccccc}a_{1}&a_{0}&0&\cdot&\cdot&\cdot&\cdot&\cdot&0\\ 0&a_{1}&a_{0}&\cdot&\cdot&\cdot&\cdot&\cdot&0\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&a_{1}&a_{0}\\ c_{0}&c_{1}&\cdot&\cdot&\cdot&\cdot&\cdot&c_{l-1}&c_{l}\end{array}\right)=\det\left(\begin{array}[]{cc}a_{1}&a_{0}\\ p&q\end{array}\right)^{l},$
where $c_{i}={l\choose i}p^{l-i}q^{i}~(i=0,1,\cdots,l)$.
The proof is straightforward and omitted.
Example 2.
Let P(z)=3z2P(z)=3z-2. Then since $\det\left(\begin{array}[]{cc}3&-2\\ -1&1\end{array}\right)=1$, $\det A=1$ holds for
$A=\left(\begin{array}[]{ccccc}3&-2&0&0&0\\ 0&3&-2&0&0\\ 0&0&3&-2&0\\ 0&0&0&3&-2\\ 1&-4&6&-4&1\end{array}\right)$
by Lemma 2. Hence, $A^{-1}$ is an integer matrix. In fact,
$A^{-1}=\left(\begin{array}[]{ccccc}-5&18&-20&8&16\\ -8&27&-30&12&24\\ -12&40&-45&18&36\\ -18&60&-68&27&54\\ -27&90&-102&40&81\end{array}\right).$
A lifting matrix is obtaind as
$\Lambda=\left(\begin{array}[]{c}16\\ 24\\ 36\\ 54\\ 81\end{array}\right)(1~-4~~6~-4~~1)=\left(\begin{array}[]{ccccc}16&-64&96&-64&16\\ 24&-96&144&-96&24\\ 36&-144&216&-144&36\\ 54&-216&324&-216&54\\ 81&-324&486&-324&81\end{array}\right)$
It holds that $\left(\cdots,\frac{14}{15},\frac{2}{5};\frac{3}{5},\frac{2}{5},\frac{1}{10},\cdots\right)\in\Omega_{P}$. In fact, $M_{P}^{4}\left(\begin{array}[]{c}\frac{14}{15}\\ \frac{2}{5}\\ \frac{3}{5}\\ \frac{2}{5}\\ \frac{1}{10}\end{array}\right)=\left(\begin{array}[]{c}2\\ 0\\ 1\\ 1\end{array}\right)$. A local lift of it is obtained by $\Lambda\left(\begin{array}[]{c}\frac{14}{15}\\ \frac{2}{5}\\ \frac{3}{5}\\ \frac{2}{5}\\ \frac{1}{10}\end{array}\right)=\left(\begin{array}[]{c}\frac{344}{15}\\ \frac{172}{5}\\ \frac{258}{5}\\ \frac{387}{5}\\ \frac{1161}{10}\end{array}\right)$. In fact,
$\left(\begin{array}[]{c}\frac{14}{15}\\ \frac{2}{5}\\ \frac{3}{5}\\ \frac{2}{5}\\ \frac{1}{10}\end{array}\right)\equiv\left(\begin{array}[]{c}\frac{344}{15}\\ \frac{172}{5}\\ \frac{258}{5}\\ \frac{387}{5}\\ \frac{1161}{10}\end{array}\right)~(\mbox{C}~\mathbb{Z})~\mbox{and}~M_{P}^{4}\left(\begin{array}[]{c}\frac{344}{15}\\ \frac{172}{5}\\ \frac{258}{5}\\ \frac{387}{5}\\ \frac{1161}{10}\end{array}\right)=\left(\begin{array}[]{c}0\\ 0\\ 0\\ 0\\ 0\end{array}\right).$
If the relation $3x-2y\equiv 0~(\rm{mod}~\mathbb{Z})$ is considered as a function $x\mapsto y$, it is a 2-valued function, and the inverse function $y\mapsto x$ is a 3-valued function. Hence, the trajectories to the forward starting at $x_{0}$ is represented by the paths on a 2-tree, and to the backward by the paths on a 3-tree. For example, the trajectory on the tee of $\left(\cdots,\frac{14}{15},\frac{2}{5};\frac{3}{5},\frac{2}{5},\frac{1}{10},\cdots\right)\in\Omega_{P}$ is as follows:
$\frac{3}{5}$$\frac{9}{10}$$\frac{2}{5}$$x_{0}\vdots$$\frac{1}{10}$$\frac{3}{5}$$\frac{2}{5}$$\frac{11}{15}$$\frac{1}{15}$$\frac{14}{15}$$\frac{3}{5}$$\frac{4}{15}\cdots\cdots\cdots\cdots$
It is also represented as a trajectory on the graph of $(x,y)\in[0,1)\times[0,1)$ satisfying 3x2y=03x-2y=0 as follows, where the numbers along the lines are $\xi$-values. In our case, $\xi=(\cdots,2;0,1,1,\cdots)$.
[…]
By the Kaneko formula, we have
$P(z)^{-1}=\left(\cdots,\frac{1}{3}\left(\frac{2}{3}\right)^{3},\frac{1}{3}\left(\frac{2}{3}\right)^{2},\frac{1}{3}\left(\frac{2}{3}\right),\frac{1}{3};~0^{\infty}\right)$
and for example,
$\frac{14}{15}=x_{-2}=2\times\frac{1}{3}+0\times\frac{1}{3}\left(\frac{2}{3}\right)+1\times\frac{1}{3}\left(\frac{2}{3}\right)^{2}+1\times\frac{1}{3}\left(\frac{2}{3}\right)^{3}+\cdots.$
The topological entropy of this system is $\log 3$, which comes from 2 factors:
(i) the choice of branches whose entropy is $\log 2$,
(ii) the local expantion whose entropy is $\log(3/2)$.
5 Factorization of polynomials and stream zeros
For a preparation, we prove the following Lemma which may be well known.
Lemma 3.
For a primitive polynomial $P(z)=a_{k}z^{k}+\cdots+a_{1}z+a_{0}\in\mathbb{Z}[x]$ with $k\geq 1,~a_{0},a_{k}\neq 0$ and $\Delta\in\mathbb{Z}_{\geq 2}$, $\Omega_{P}\cap\{0,\frac{1}{\Delta},\frac{2}{\Delta},\cdots,\frac{\Delta-1}{\Delta}\}^{\mathbb{Z}}$ is a finite set.
Proof.
Clearly our lemma follows from the following Fact 1, and Fact 1 follows from Fact 2.
Fact 1: If $X=(x_{n};n\in\mathbb{Z})\in St(\mathbb{Z})$ satisfies that $P(z)\times X=(0^{\infty};0^{\infty})~({\rm mod}~\Delta)$, then $X~({\rm mod}~\Delta)$ is periodic with the period bounded by a function of $k$ and $\Delta$.
Fact 2: Let $q$ be a power of a prime. If $X=(x_{n};n\in\mathbb{Z})\in St(\mathbb{Z})$ satisfies that $P(z)\times X=(0^{\infty};0^{\infty})~({\rm mod}~q)$, then $X~({\rm mod}~q)$ is periodic with the period bounded by a function of $k$ and $q$.
To prove Fact 2, we need the following Fact 3.
Fact 3: Let $p$ be a prime and $B=(b_{n};n\in\mathbb{Z})\in St(\mathbb{Z})$ be periodic. If $X=(x_{n};n\in\mathbb{Z})\in St(\mathbb{Z})$ satisfies that $P(z)\times X=B~({\rm mod}~p)$. Then, $X~({\rm mod}~p)$ is periodic with the period bounded by a function of $k$ and the period of $B$.
Proof.
Let $K$ be the period of $B$. Let $h$ be the minimum such that $a_{h}\neq 0~({\rm mod}~p)$, which exists since $P(z)$ is primitive. Then for any $n\in\mathbb{Z}$, given $(x_{n-k+h},\cdots,x_{n-1})~({\rm mod}~p)$, $x_{n}~({\rm mod}~p)$ is determined so that
$a_{k}x_{n-k+h}+\cdots+a_{h}x_{n}+\cdots+a_{0}x_{n+h}=b_{n+h}~({\rm mod}~p)$
for any $n\in\mathbb{Z}$. That is, $x_{n}$ is determined as above by
$(x_{n-k+h},\cdots,x_{n-1},i)\in\{0,1,\cdots,p-1\}^{k-h}\times\{0,1,\cdots,K-1\},$
where $i$ corresponds to the $i$-th element of a cycle of $B$. Hence, there exists $n<m$ with $m-n\leq p^{k-h}K$ such that $x_{n}$ and $x_{m}$ correspond to the same element in the above, and hence, $x_{n}=x_{m}$. This implies that $x_{n+1}=x_{m+1},~x_{n+2}=x_{m+2},$$\cdots$. Thus, $X$ is periodic with the period at most $p^{k-h}K$, which prove Fact 3.
Now, we prove Fact 2 for $q=p^{l}$ by the induction on $l=1,2,\cdots$, where $p$ is an arbitrary prime. If $l=1$, then Fact 2 is nothing but Fact 3 with $B=(0^{\infty};0^{\infty})$. Assume that Fact 2 holds for $l=l_{0}\geq 1$. Let $X=(x_{n};n\in\mathbb{Z})\in St(\mathbb{Z})$ satisfy that $P(z)\times X=(0^{\infty};0^{\infty})~({\rm mod}~p^{l_{0}+1})$. Let $x_{n}=x_{n}^{\prime}+x_{n}^{\prime\prime}p^{l_{0}}$ with $x_{n}^{\prime}\in\{0,1,\cdots,p^{l_{0}}-1\}$ and $x_{n}^{\prime\prime}\in\mathbb{Z}$ for any $n\in\mathbb{Z}$. Since $P(z)\times X=(0^{\infty};0^{\infty})~({\rm mod}~p^{l_{0}+1})$, $P(z)\times(x_{n}^{\prime};n\in\mathbb{Z})=(0^{\infty};0^{\infty})~({\rm mod}~p^{l_{0}})$. Hence, by the assumption of the induction $(x_{n}^{\prime};n\in\mathbb{Z})$ is periodic with the period bounded by a function of $k$ and $p^{l_{0}}$. Let $b_{n}p^{l_{0}}=a_{k}x^{\prime}_{n-k}+\cdots+a_{1}x^{\prime}_{n-1}+a_{0}x^{\prime}_{n}~(\forall n\in\mathbb{Z})$. Then, $B=(b_{n};n\in\mathbb{Z})\in St(\mathbb{Z})$ is periodic with the period which is a factor of that of $(x^{\prime}_{n};n\in\mathbb{Z})$, hence, a factor of $p^{l_{0}}$. Then we have
$a_{k}x_{n-k}+\cdots+a_{1}x_{n-1}+a_{0}x_{n}\\ =(a_{k}x^{\prime\prime}_{n-k}+\cdots+a_{1}x^{\prime\prime}_{n-1}+a_{0}x^{\prime\prime}_{n}+b_{n})p^{l_{0}}=0~~({\rm mod}~p^{l_{0}+1}).$
Hence,
$a_{k}x^{\prime\prime}_{n-k}+\cdots+a_{1}x^{\prime\prime}_{n-1}+a_{0}x^{\prime\prime}_{n}=-b_{n}~~({\rm mod}~p).$
Thererfor by Fact 3, $(x_{n}^{\prime\prime};n\in\mathbb{Z})$ is periodic with the period bounded by a function of $k$ and $p^{l_{0}}$. This implies that $X~({\rm mod}~p^{l_{0}+1})$ is periodic with a period bounded by a function of $k$ and $p^{l_{0}+1}$, which is Fact 2 for $l=l_{0}+1$. Thus, Fact 2 holds. ∎
For two primitive polynomials $P(z),Q(z)\in\mathbb{Z}[x]$ with degree $k$ and $h$, respectively.
$\begin{array}[]{cc}P(z)=a_{k}z^{k}+\cdots+a_{1}z+a_{0}&(a_{k}\neq 0,a_{0}\neq 0)\\ Q(z)=b_{h}z^{h}+\cdots+b_{1}z+b_{0}&(b_{h}\neq 0,b_{0}\neq 0).\end{array}$ (6)
We define the resultant matrix $\mathcal{R}$ between them which is a square matrix of degree $k+h$ as follows:
$\mathcal{R}=\left(\begin{array}[]{cccccc}a_{k}&a_{k-1}&\cdots&a_{0}&0&\cdots\\ \cdots&\ddots&\ddots&\ddots&\ddots&\ddots\\ \cdots&0&a_{k}&a_{k-1}&\cdots&a_{0}\\ b_{h}&b_{h-1}&\cdots&b_{0}&0&\cdots\\ \cdots&\ddots&\ddots&\ddots&\ddots&\ddots\\ \cdots&0&b_{h}&b_{h-1}&\cdots&b_{0}\end{array}\right).$
Then, it is well known [10] that $P(z),Q(z)$ are coprime if and only if $\mathcal{R}$ is regular. It also holds that if $P(z),Q(z)$ are coprime, then there exists polynomials $A(z),B(z)$ with integer coefficients such that
$\deg A<h,~\deg B<k~\mbox{and}~A(z)\times P(z)+B(z)\times Q(z)=\det\mathcal{R}\in\mathbb{Z}\setminus\{0\}.$
Denote $\det(\mathcal{R})$ by $Res(P,Q)$. For a stream zero’s of $P(z)$, define $\dim\Omega_{P}$ to be the maximal $l$ such that
$\{(x_{0},x_{1},\cdots,x_{l-1});~(x_{n},n\in\mathbb{Z})\in\Omega_{P}\}=\mathbb{T}^{l}.$
Clearly, $\dim\Omega_{P}=k,~\dim\Omega_{Q}=h$ for $P,Q$ in (6).
For $x=(x_{n};n\in\mathbb{Z})$ and $y=(y_{n};b\in\mathbb{Z})\in St(\mathbb{R})$, define the Hadamard product
$x\odot y:=(x_{n}y_{n};n\in\mathbb{Z}).$
Let $\mathbf{n}=(\cdots,-1;0,1,2,\cdots)\in St(\mathbb{Z})$ and $i=1,2,\cdots$, we define
$\mathbf{n}^{\odot i}=\underbrace{\mathbf{n}\odot\cdots\odot\mathbf{n}}_{\mbox{i times}}=(n^{i};n\in\mathbb{Z})\in St(\mathbb{Z}).$
We now show there are close relationships between the factorization of a primitive polynomial $P(z)\in\mathbb{Z}[z]$ and a decomposition of $\Omega_{P}$ into the corresponding summands.
Theorem 4.
[…]
(i) $P(z)$ and $Q(z)$ are coprime if and only if $Res(P,Q)\neq 0$. In this case, we have
$\Omega_{P}\cap\Omega_{Q}\subset\{\frac{0}{\Delta},\frac{1}{\Delta},\cdots,\frac{\Delta-1}{\Delta}\}^{\mathbb{Z}}$
with $\Delta=Res(P,Q)$. Therefore, $P(z)$ and $Q(z)$ are coprime if and only if $\Omega_{P}\cap\Omega_{Q}$ is a finite set.
(ii) $P(z)$ is a factor of $Q(z)$ if and only if $\Omega_{P}\subset\Omega_{Q}$.
(iii) If $P(z),Q(z)$ are coprime, then we have $\Omega_{P\times Q}=\Omega_{P}\oplus\Omega_{Q}$, where $\oplus$” implies the semi-direct sum, that is, the intersection of the summands is a finite set.
(iv) If all the roots of $P(z)=0$ are simple, then for any $H=1,2,\cdots$, we have
$\Omega_{P^{H}}=\sum_{i=0}^{H-1}\mathbf{n}^{\odot i}\odot\Omega_{P}.$
Proof.
(i) Let $P(z),Q(z)$ be coprime. Then, the resultant matrix $\mathcal{R}$ is regular and the entries of $\mathcal{R}^{-1}$ are in $\mathbb{Z}/\Delta$, where $\Delta=Res(P,Q)\in\mathbb{Z}\setminus\{0\}$. Since for any $k\in\mathbb{Z}$ and $\xi=(\xi_{i})_{i\in\mathbb{Z}}\in\Omega_{P}\cap\Omega_{Q}$, we have $\mathcal{R}\left(\begin{array}[]{c}\xi_{k+1}\\ \xi_{k+2}\\ \cdots\\ \xi_{k+n+m}\end{array}\right)\in\mathbb{Z}^{k+n+m}$, it holds that $\left(\begin{array}[]{c}\xi_{k+1}\\ \xi_{k+2}\\ \cdots\\ \xi_{n+m}\end{array}\right)\in\frac{1}{\Delta}\mathbb{Z}^{n+m}$. This implies that $\Omega_{P}\cap\Omega_{Q}\subset\{\frac{0}{\Delta},\frac{1}{\Delta},\cdots,\frac{\Delta-1}{\Delta}\}^{\mathbb{Z}}$, and hence, $\Omega_{P}\cap\Omega_{Q}$ ia a finite set by Lemma 3.
Conversely, if $P(z),Q(z)$ are not coprime, then there exists a common factor of them with degree at least 1. Hence we have $\dim(\Omega_{P}\cap\Omega_{Q})\geq 1$, and hence, $\Omega_{P}\cap\Omega_{Q}$ is an infinite set.
(ii) Assume that $P(z)$ is a factor of $Q(z)$ so that $Q(z)=P(z)\times R(z)$, where $R(z)$ is a polynomial with integer coefficients. Then for any $\xi\in\Omega_{P}$, we have
$Q(z)\times\xi=(P(z)\times R(z))\times\xi=R(z)\times(P(z)\times\xi)=(0^{\infty};0^{\infty}).$
Hence, we have $\xi\in\Omega_{Q}$ and $\Omega_{P}\subset\Omega_{Q}$.
Conversely, if $P(z)$ is not a factor of $Q(z)$, then there exists a nontrivial factor $R(z)$ of $P(z)$ such that $R(z),Q(z)$ are coprime. Then by (i), $\Omega_{Q}\cap\Omega_{R}$ is a finite set, while $\Omega_{R}$ is not. Hence, $\Omega_{R}\subset\Omega_{Q}$ does not hold. Thus, $\Omega_{P}\subset\Omega_{Q}$ does not hold since $\Omega_{P}\supset\Omega_{R}$.
(iii) By (ii) and the fact that $\Omega_{P\times Q}$ is a additive group, it holds that
$\Omega_{P\times Q}\supset\Omega_{P}+\Omega_{Q}.$ (7)
Since $P(z),Q(z)$ are coprime, there exist polynomials with integer coefficients $A(z),B(z)$ with $\deg A(z)\leq h$ and $\deg B(z)\leq k$ such that
$P\times A+Q\times B=\Delta,$ (8)
where $\Delta=Res(P,Q)$. By (8), we have
$P\times Q\times B=P\times(\Delta-P\times A),~P\times Q\times A=Q\times(\Delta-Q\times B).$ (9)
Take $X:=(x_{n};n\in\mathbb{Z})\in\Omega_{P\times Q}$ arbitrary. By Corollary 1, there exists $Y:=(y_{n};n\in\mathbb{Z})\in\Omega_{P\times Q}$ such that $\Delta Y=X$. Since $Y\in\Omega_{P\times Q}$, we have by (9) that
$P\times(\Delta-P\times A)\times Y=Q\times(\Delta-Q\times B)\times Y=(0^{\infty};0^{\infty}).$
Let
$(u_{n};n\in\mathbb{Z})=(\Delta-P\times A)\times Y,~(v_{n};n\in\mathbb{Z})=(\Delta-Q\times B)\times Y.$
Then, from above $(u_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $(v_{n};n\in\mathbb{Z})\in\Omega_{Q}$. Moreover, by (8),
$(u_{n};n\in\mathbb{Z})+(v_{n};n\in\mathbb{Z})=(2\Delta-P\times A-Q\times B)\times Y=\Delta Y=X.$
Thus, we have $\Omega_{P\times Q}\subset\Omega_{P}+\Omega_{Q}$, hence together with (7) and (i), we complete the proof.
(iv) Considering the lifts, the equation
$\Omega_{P^{H}}=\sum_{i=0}^{H-1}\mathbf{n}^{\odot i}\odot\Omega_{P}$
follows from the equation
$\tilde{\Omega}_{P^{H}}=\sum_{i=0}^{H-1}\mathbf{n}^{\odot i}\odot\tilde{\Omega}_{P}.$
For any $X=(x_{n};n\in\mathbb{Z})\in St(\mathbb{T})$, define the derivative $X^{\prime}\in St(\mathbb{R})$ to be $\sum_{n\in\mathbb{Z}}nx_{n}z^{n-1}$ as usual. That is, $X^{\prime}=\sigma(\mathbf{n}\odot X)$, where $\sigma$ is the shift on $St(\mathbb{T})$. Clearly, if $X\in\Omega_{P}$, then $(P\times X)^{\prime}=(0^{\infty};0^{\infty})$.
Hence if $X\in\Omega_{P^{H-1}}$, then we have
$(0^{\infty};0^{\infty})=(P^{H}\times X)^{\prime}=HP^{H-1}P^{\prime}\times X+P^{H}\times X^{\prime}=P^{H}\times\sigma(\mathbf{n}\odot X).$
Therefore, $\sigma(\mathbf{n}\odot X)\in\Omega_{P^{H}}$ and $\mathbf{n}\odot X\in\Omega_{P^{H}}$. Thus, $\mathbf{n}\odot\Omega_{P^{H-1}}\subset\Omega_{P^{H}}$. It follows from this that
$\Omega_{P^{H}}\supset\sum_{i=0}^{H-1}\mathbf{n}^{\odot i}\odot\Omega_{P}~~(H=1,2,\cdots).$ (10)
To prove the oposite inclusion, we use the local lifts of these terms. By Corollary 1,
$\tilde{\Omega}_{P^{H}}|_{N}\supset\sum_{i=0}^{H-1}\mathbf{n}^{\odot i}\odot\tilde{\Omega}_{P}|_{N}$
holds for any large $N\in\mathbb{Z}_{>0}$. Since the both sides are vector spaces over $\mathbb{R}$, they coincide if both sides have the same dimension $kH$ as that of the vector spaces over $\mathbb{R}$.
For this purpose, we need the following lemma which is a generalization of the Vandermonde’s determinant due to R. P. Flowe and A. G. Harris [4].
Lemma 4.
[4] Let $l_{1},\cdots,l_{h}$ and $L$ be positive integers such that $l_{1}+\cdots+l_{h}=L$. Let
$\mathcal{M}[x_{1},\cdots,x_{h};L;l_{1},\cdots,l_{h}]:=\left(\mathcal{M}[x_{1},L,l_{1}]\mathcal{M}[x_{2},L,l_{2}]\cdots\mathcal{M}[x_{h},L,l_{h}]\right)$
be the $L\times L$-matrix whose columns from 1st to $l_{1}$-th is $\mathcal{M}[x_{1},L,l_{1}]$ and from $l_{1}+1$-th to $l_{1}+l_{2}$-th is $\mathcal{M}[x_{2},L,l_{2}]$, $\cdots$, where $x_{1},x_{2},\cdots,x_{h}\in\mathbb{C}$ and
$\mathcal{M}(x,L,l)=\left(\begin{array}[]{ccccc}1&0&0&\cdots&0\\ x&x&x&\cdots&x\\ x^{2}&2x^{2}&2^{2}x^{2}&\cdots&2^{l-1}x^{2}\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ \cdot&\cdot&\cdot&\cdots&\cdot\\ x^{L-1}&(L-1)x^{L-1}&(L-1)^{2}x^{L-1}&\cdots&(L-1)^{l-1}x^{L-1}\end{array}\right).$
Then
$\det\mathcal{M}[x_{1},\cdots,x_{h};L;l_{1},\cdots,l_{h}]=C\prod_{i=1}^{h}x_{i}^{l_{i}(l_{i}-1)/2}\prod_{i<j}(x_{j}-x_{i})^{l_{i}l_{j}}$ (11)
with $C=\prod_{i=1}^{h}\prod_{j=1}^{l_{i}}j!\neq 0$.
Continuation of the proof of (iv) of Theorem 4:  Let $x_{1},x_{2},\cdots,x_{s}$ be distinct real numbers and
$y_{1},y_{2},\cdots,y_{u};\overline{y}_{1},\overline{y}_{2},\cdots,\overline{y}_{u}$
be distinct complex numbers not being real which are the set of roots of $P(z)=0$, where $\overline{y}$ is the complex conjugate of $y$. Then, $s+2u=k$ and the roots of $P(z)^{H}=0$ are as above with the multiplicity $H$. Hence, by Lemma 2, the set of vectors
$\{\mathbf{n}^{\odot i}(x_{i}^{n};n\in\mathbb{Z});~i=0,1,\cdots,H-1;j=1,2,\cdots,s\}$
$\cup$ $\{\mathbf{n}^{\odot i}(y_{i}^{n};n\in\mathbb{Z});~i=0,1,\cdots,H-1;j=1,2,\cdots,u\}$
$\cup$ $\{\mathbf{n}^{\odot i}(\overline{y}_{i}^{n};n\in\mathbb{Z});~i=0,1,\cdots,H-1;j=1,2,\cdots,u\}$
spans $kH$-dimensional vector space over $\mathbb{C}$. Replacing the pair
$(y_{i}^{n};n\in\mathbb{Z}),~(\overline{y}_{i}^{n};n\in\mathbb{Z})~\mbox{by}~(\Re(y_{i}^{n});n\in\mathbb{Z}),~(\Im(y_{i}^{n});n\in\mathbb{Z}),$
we get the same space whose restriction to the real parts is spanned by the replaced one. This vector space over $\mathbb{R}$ is same as the vector space spanned by
$\tilde{\Omega}_{P},~\mathbf{n}\odot\tilde{\Omega}_{P},~\cdots,~\mathbf{n}^{\odot(H-1)}\tilde{\Omega}_{P},$
and hence, has the dimension $kH$ which coincides the dimension of the space $\tilde{\Omega}_{P^{H}}$. Thus, we proved the oposite inclusion of (10). ∎
The following Theorem is already known in S. Akiyama, T. Kamae & H. Kaneko [8] except for the statement of semi-direct.
Theorem 5.
Let $P(z)\in\mathbb{Z}[z]$ be a primitive polynomial with the resolution into irreducible factors
$P(z)=Q_{1}(z)^{r_{1}}\cdots Q_{h}(z)^{r_{h}},$
where $Q_{i}(z)\in\mathbb{Z}[z]~(i,1,\cdots,h)$. Then, we have
$\Omega_{P}=\sum_{i=1}^{k}\sum_{j=1}^{r_{i}-1}\mathbf{n}^{\odot j}\odot\Omega_{Q_{j}},$
where the sum is semi-direct.
Proof.
By (iii) and (iv) of Theorem 4, we have
$\Omega_{P}=\sum_{i=1}^{k}\Omega_{Q_{i}}^{r_{i}}$
$=\sum_{i=1}^{k}\sum_{j=1}^{r_{i}-1}\mathbf{n}^{\odot j}\odot\Omega_{Q_{j}}.$
As a consequence of Theorem 5, we have the following Corollary.
Corollary 2.
For polynomials $P(z),Q(z),R(z)$ with integer coefficients, assume that $P(z)=Q(z)\times R(z)$ and $Res(Q,R)=\pm 1$. Then dynamical systems $(\Omega_{P},\sigma)$ and $(\Omega_{Q},\sigma)\otimes(\Omega_{R},\sigma)$ are conjugate, where $\otimes$” implies the direct product of dynamical systems. Hence, dynamical systems $(\Xi_{P},\sigma)$ and $(\Xi_{Q},\sigma)\otimes(\Xi_{R},\sigma)$ are conjugate.
6 Stream zeros and the strong automorphism group
Let
$P(z)=a_{k}z^{k}+a_{k-1}z^{k-1}+\cdots+a_{1}z+a_{0}\in\mathbb{Z}[z]~(k\geq 1,a_{0}\neq 0,a_{k}\neq 0)$ (12)
be a primitive polynomial and
$C_{P}=\left(\begin{array}[]{ccccc}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ \cdots&\cdots&\cdots&\ddots&\cdots\\ 0&0&\cdots&\cdots&1\\ -\frac{a_{k}}{a_{0}}&-\frac{a_{k-1}}{a_{0}}&\cdots&\cdots&-\frac{a_{1}}{a_{0}}\end{array}\right).$ (13)
be the companion matrix.
Since $C_{P}$ is not an integer matrix if $a_{0}\neq\pm 1$, the action of $C_{P}$ to $\mathbb{T}^{k}$ is multivalued. In fact, we generally define the action of a $(k\times k)$-matrix $F$ with rationl entries to $\mathbb{T}^{k}$ as $\pi\circ F\circ\pi^{-1}$, where $\pi:\mathbb{R}^{k}\to\mathbb{T}^{k}$ is the coordinatewise projection. We call $\pi\circ F\circ\pi^{-1}$ the induced action of $F$ and denote $\overline{F}$. For example, if $F=(p/q)I_{1}$ ($I_{k}$ being the identity matrix of degree $k$) with irreducible fraction $p/q$ and $x\in\mathbb{T}$, then
$\overline{F}x:=\pi\left(\frac{p}{q}\times\{\{x\}+i;~i\in\mathbb{Z}\}\right)=\left\{\frac{\{px\}}{q}+\frac{i}{q};~i=0,1,\cdots,q-1\right\},$
where $\{x\}$ being the representative of $x\in\mathbb{T}$ in $[0,1)$. Hence,
$\Omega_{P}=$ $\{(x_{n};n\in\mathbb{Z})\in\mathbb{T}^{\mathbb{Z}}:~a_{k}x_{n-k}+\cdots+a_{1}x_{n-1}+a_{0}x_{n}=0~(\forall n\in\mathbb{Z})\}$
$=$ $\left\{(x_{n};n\in\mathbb{Z})\in St(\mathbb{T});~\left(\begin{array}[]{c}x_{n-k}\\ \\ \cdot\\ \cdot\\ x_{n-1}\\ x_{n}\end{array}\right)\in\overline{C}_{P}\left(\begin{array}[]{c}x_{n-k-1}\\ \cdot\\ \cdot\\ x_{n-2}\\ x_{n-1}\end{array}\right)~(\forall n\in\mathcal{N})\right\}$
holds.
Lemma 5.
(i) The correspondence $F\mapsto\overline{F}$ defined on the set of $(k\times k)$-matrices with rational entries is one-to-one.
(ii) Any $(k\times k)$-integral matrix $F$ can be identified with $\overline{F}$ consistently. Moreover, for any $(k\times k)$-matrix $G$ with rational entries, we have
(1) $\overline{FG}=F\overline{G}$
(2) $\overline{GF}=\overline{G}F$.
Proof.
(i) Assume that $F\neq G$. Then, there exists $v\in\mathbb{R}^{k}$ such that (FG)v(F-G)v is a nonzero irrational vector. Let $u=\pi v$, Then, there exist $S,S^{\prime}\subset\mathbb{Q}^{k}$ such that $\overline{F}u=\pi F\pi^{-1}u=\pi(Fv+S)$ and $\overline{G}u=\pi G\pi^{-1}u=\pi(Gv+S^{\prime})$. Therefore, $\mathbb{O}\neq\pi(F-G)v\in(\overline{F}-\overline{G})u$. Hence, $\overline{F}\neq\overline{G}$.
(ii) Since $Fw$ is an integer vector if $w$ is so, for any $v\in\pi^{-1}u$ with $u\in\mathbb{T}^{k}$, $Fv\equiv F\{u\}~(\mbox{mod}~\mathbb{Z}^{k})$. That is, $\overline{F}u=\pi F\pi^{-1}u\equiv F\{u\}~(\mbox{mod}~\mathbb{Z}^{k})$. In this sense, we can identify $\overline{F}$ with $F$.
(1) Note that $\pi F\pi^{-1}\pi=\pi F$. Hence,
$\overline{FG}=\pi FG\pi^{-1}=\pi F\pi^{-1}\pi G\pi^{-1}=\overline{F}~\overline{G}=F\overline{G}.$
(2) Take an arbitrary $u\in\mathbb{T}^{k}$ and let $\pi v=u$ for some $v\in\mathbb{R}^{k}$. Since $F\mathbb{Z}^{k}=\mathbb{Z}^{k}$, it holds that
$\overline{GF}u=\pi GF\pi^{-1}u=\pi GF(v+\mathbb{Z}^{k})=\pi G(Fv+\mathbb{Z}^{k})=\pi G\pi^{-1}Fu=\overline{G}Fu.$
Thus, $\overline{GF}=\overline{G}F$ holds. ∎
Let $\Omega_{P}$ be the set of stream zero’s of $P(z)$. A mapping $\varphi:\Omega_{P}\to\Omega_{P}$ is called an automorphism of $\Omega_{P}$ if it is an isomorphism of $\Omega_{P}$ as the topological linear space over $\mathbb{Z}$ such that $\varphi\circ\sigma=\sigma\circ\varphi$, where $\sigma$ is the shift. Furthermore, we call it a strong automorphism if for any $(x_{n};n\in\mathbb{Z}),~(y_{n};n\in\mathbb{Z})\in\Omega_{P}$ with $\varphi(x_{n};n\in\mathbb{Z})=(y_{n};n\in\mathbb{Z})$,
$(y_{0},y_{1},\cdots,y_{k-1})\mbox{ is determined by }(x_{0},x_{1},\cdots,x_{k-1}),~\mbox{and}$
$(x_{0},x_{1},\cdots,x_{k-1})\mbox{ is determined by }(y_{0},y_{1},\cdots,y_{k-1}).$
Since a strong atomorphis $\varphi$ commutes with the shift, it holds for any $n\in\mathbb{Z}$ that
$(y_{n},y_{n+1},\cdots,y_{n+k-1})\mbox{ is determined by }(x_{n},x_{n+1},\cdots,x_{n+k-1}),~\mbox{and}$
$(x_{n},x_{n+1},\cdots,x_{n+k-1})\mbox{ is determined by }(y_{n},y_{n+1},\cdots,y_{n+k-1}).$
Note that the shift $\sigma$ on $\Omega_{P}$ is an automorphism, it is a strong automorphism as well only if $a_{0}=\pm 1$ and $a_{k}=\pm 1$.
Definition 4.
By $Saut_{P}$, we denote the set of $k\times k$ integer matrices B such that $BC_{P}=C_{P}B$ and $\det B=\pm 1$.
Note that $BC_{P}^{-1}=C_{P}^{-1}B$ also follows if $B\in Saut_{P}$ since
$BC_{P}^{-1}=C_{P}^{-1}C_{P}BC_{P}^{-1}=C_{P}^{-1}BC_{P}C_{P}^{-1}=C_{P}^{-1}B.$
The following theorem states that each of strong automorphisms on $\Omega_{P}$ can be represented by a matrix from $Saut_{P}$.
Theorem 6.
[…]
(i) For any $B\in Saut_{P}$, let $B=(b_{i,j})_{1\leq i,,j\leq k}$. Then, $b_{1,k},b_{2,k}\cdots,b_{k-1,k}$ are multiples of $a_{0}$, while $b_{k.k}$ and $a_{0}$ are coprime.
(ii) A mapping $\varphi:\Omega_{P}\to\Omega_{P}$ is a strong automorphism if and only there exists $B\in Saut_{P}$ such that for any $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $n\in\mathbb{Z}$,
$\left(\begin{array}[]{c}y_{n}\\ y_{n+1}\\ \cdot\\ \cdot\\ y_{n+k-1}\end{array}\right)=B\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\\ \cdot\\ \cdot\\ x_{n+k-1}\end{array}\right).$ (14)
holds, where $(y_{n};n\in\mathbb{Z}):=\varphi(x_{n};n\in\mathbb{Z})$. Therefore, a strong automorphism $\varphi$ can be identified with $B\in Saut_{P}$ as above.
Proof.
(i) Assume that $a_{0}\neq\pm 1$ since otherwise there is nothing to prove. Since $BC_{P}=C_{P}B$, by comparing their 1st raws, we have
$\frac{a_{k}}{a_{0}}b_{1,k}=b_{2,1},~b_{1,1}-\frac{a_{k-1}}{a_{0}}b_{1,k}=b_{2,2},~\cdots,~b_{1,k-1}-\frac{a_{1}}{a_{0}}b_{1,k}=b_{2,k}.$
As $B$ is an integer matrix, it follows that $\frac{a_{k}}{a_{0}}b_{1,k},\frac{a_{k-1}}{a_{0}}b_{1,k},\cdots,\frac{a_{1}}{a_{0}}b_{1,k}$ are integers. Since $P(z)$ is primitive and GCD of $a_{0},a_{1},\cdots,a_{k}$ is 1, this is possible only if $b_{1,k}$ is a multiple of $a_{0}$. In the same way, by comparing the 2nd, 3rd, … , (k1)(k-1)-th rows, it holds that $b_{1,k},b_{2,k},\cdots,b_{k-1,k}$ are multiples of $a_{0}$. Suppose that $b_{k.k}$ and $a_{0}$ are not coprime. Then, $b_{1,k},b_{2,k},\cdots,b_{k-1,k}$ and $b_{k,k}$ have common factor larger than 1. This contradicts with $\det B=\pm 1$.
(ii) Let $\varphi:\Omega_{P}\to\Omega_{P}$ be a strong automorphism. For any $X\in\Omega_{P}$ with $X=(x_{n};n\in\mathbb{Z})$, let $Y:=(y_{n};~n\in\mathbb{Z})=\varphi(x_{n};n\in\mathbb{Z})\in\Omega_{P}$. Then, for any $n\in\mathbb{Z}$,
$(y_{n},y_{n+1},\cdots,y_{n+k-1})\mbox{ is determined by }(x_{n},x_{n+1},\cdots,x_{n+k-1}),\mbox{ and}$
$(x_{n},x_{n+1},\cdots,x_{n+k-1})\mbox{ is determined by }(y_{n},y_{n+1},\cdots,y_{n+k-1}).$
Since $\varphi$ is an automorphism and $\Omega|_{\{n,n+1,\cdots,n+k-1\}}=\mathbb{T}^{k}~(\forall n\in\mathbb{Z})$, there exist integer matrices $\cdots,B_{-1},B_{0},B_{1},\cdots$ whose inverses are also integer matrices such that
$\left(\begin{array}[]{c}y_{n}\\ y_{n+1}\\ \cdot\\ \cdot\\ y_{n+k-1}\end{array}\right)=B_{n}\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\\ \cdot\\ \cdot\\ x_{n+k-1}\end{array}\right)$
holds for any $n\in\mathbb{Z}$. Since $\varphi\circ\sigma=\sigma\circ\varphi$, we have
$B_{1}\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \cdot\\ \cdot\\ x_{k}\end{array}\right)=B_{1}\left(\begin{array}[]{c}(\sigma X)_{0}\\ (\sigma X)_{1}\\ \cdot\\ \cdot\\ (\sigma X)_{k-1}\end{array}\right)=B_{1}B_{0}^{-1}\left(\begin{array}[]{c}(\sigma Y)_{0}\\ (\sigma Y)_{1}\\ \cdot\\ \cdot\\ (\sigma Y)_{k-1}\end{array}\right)\\ =B_{1}B_{0}^{-1}\left(\begin{array}[]{c}y_{1}\\ y_{2}\\ \cdot\\ \cdot\\ y_{k}\end{array}\right)=B_{1}B_{0}^{-1}B_{1}\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \cdot\\ \cdot\\ x_{k}\end{array}\right).$
Therefore we have $B_{1}=B_{1}B_{0}^{-1}B_{1}$, and hence, $B_{1}=B_{0}$ follows. In the same way, it follows that $B_{n}=B_{0}$ for all $n\in\mathbb{Z}$. Let $B_{n}=B~(\forall n\in\mathbb{Z})$.
By (ii) of Lemma 4, it holds that
$\pi BC_{P}\pi^{-1}\left(\begin{array}[]{c}x_{0}\\ x_{1}\\ \cdot\\ \cdot\\ x_{k-1}\end{array}\right)=B\pi C_{P}\pi^{-1}\left(\begin{array}[]{c}x_{0}\\ x_{1}\\ \cdot\\ \cdot\\ x_{k-1}\end{array}\right)$
$=$ $B\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \cdot\\ \cdot\\ x_{k}+\{0,\frac{1}{a_{0}}\cdots\frac{a_{0}-1}{a_{0}}\}\end{array}\right)=\left(\begin{array}[]{c}y_{1}\\ y_{2}\\ \cdot\\ \cdot\\ y_{k}+\{0,\frac{1}{a_{0}}\cdots\frac{a_{0}-1}{a_{0}}\}\end{array}\right)$
$=$ $\pi C_{P}\pi^{-1}\left(\begin{array}[]{c}y_{0}\\ y_{1}\\ \cdot\\ \cdot\\ y_{k-1}\end{array}\right)=\pi C_{P}B\pi^{-1}\left(\begin{array}[]{c}x_{0}\\ x_{1}\\ \cdot\\ \cdot\\ x_{k-1}\end{array}\right)$
for any $\left(\begin{array}[]{c}x_{0}\\ x_{1}\\ \cdot\\ \cdot\\ x_{k-1}\end{array}\right)\in\mathbb{T}^{k}$, where we used (i) to show the equality between the 3rd term and the 4th term.
Hence, $\overline{BC_{P}}=\overline{C_{P}B}$ holds, from which $BC_{P}=C_{P}B$ follows by (i) of Lemma 4. Thus, $B\in Saut_{P}$.
Conversely, for $B\in Saut_{P}$, let a mapping $\varphi:\Omega_{P}\to\Omega_{P}$ satisfy (14) for any $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $(y_{n};~n\in\mathbb{Z})=\varphi(x_{n};n\in\mathbb{Z})\in\Omega_{P}$. Then, $(y_{n},y_{n+1},\cdots,y_{n+k-1})$ is determined by $(x_{n},x_{n+1},\cdots,x_{n+k-1})$. Also considering $B^{-1}$, $(x_{n},x_{n+1},\cdots,x_{n+k-1})$ is determined by $(y_{n},y_{n+1},\cdots,y_{n+k-1})$. On the other hand, the relation (14) which holds independently of $n\in\mathbb{Z}$ implies $\varphi\circ\sigma=\sigma\circ\varphi$. Thus, $\varphi$ is a strong automorphism. ∎
Using Theorem 6, we turn to investigate the structure of $Saut_{P}$.
Theorem 7.
[…]
(i) For a $k\times k$ matrix $B$, $B\in Saut_{P}$ if and only if there exists a unit $\lambda_{1}$ in $\mathbb{Q}(\theta_{1})$ with
$\lambda_{1}\in\bigcap_{i=1}^{k}(\mathbb{Z}\theta_{1}^{i-1}+\mathbb{Z}\theta_{1}^{i-2}+\cdots+\mathbb{Z}\theta_{1}^{i-k})$ (15)
such that
$B=\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}^{2}&\cdots&\theta_{k}^{2}\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)\left(\begin{array}[]{ccc}\lambda_{1}&\cdots&0\\ &\ddots&\\ 0&\cdots&\lambda_{k}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{k}&\cdots&\theta_{k}^{k}\\ \cdots&\cdots&\cdots\\ \theta_{1}^{2}&\cdots&\theta_{k}^{2}\\ \theta_{1}&\cdots&\theta_{k}\end{array}\right)^{-1}.$ (16)
[…]
(ii) $Saut_{P}$ is isomorphic to a direct product of $l$ number of cyclic groups of infinite orders with a finite group, where $l\leq k_{1}+k_{2}-1$ and $k_{1},2k_{2}$ are the numbers of real roots and complex roots of $P(z)=0$, respectively.
Proof.
(i) Assume $B\in Saut_{P}$. Since
$C_{P}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=\left(\begin{array}[]{c}\theta_{i}^{k-1}\\ \vdots\\ \theta_{i}\\ 1\end{array}\right)=\theta_{i}^{-1}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)~(i=1,2,\cdots,k)$
and $BC_{P}=C_{P}B$, we have
$C_{P}B\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=BC_{P}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=\theta_{i}^{-1}B\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)~(i=1,2,\cdots,k).$
Since the eigenspace of $C_{P}$ corresponding to each eigenvalue $\theta_{i}~(i=1,\cdots,k)$ is 1-dimensional, there exists a unique $\lambda_{i}\in\mathbb{C}$ such that
$B\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)=\lambda_{i}\left(\begin{array}[]{c}\theta_{i}^{k}\\ \vdots\\ \theta_{i}^{2}\\ \theta_{i}\end{array}\right)~(i=1,2,\cdots,k).$
Thus, we have (16). Since $B=(b_{ij})_{i,j=1,\cdots,k}$ is an integer matrix and
$b_{i1}\theta_{1}^{k}+b_{i2}\theta_{1}^{k-1}+\cdots+b_{ik}\theta_{1}=\lambda_{1}\theta^{k-i+1}$
holds for any $i=1,\cdots,k$, we have $\lambda_{1}\in\mathbb{Z}\theta_{1}^{i-1}+\mathbb{Z}\theta_{1}^{i-2}+\cdots+\mathbb{Z}\theta_{1}^{i-k}$, and hence (15). From (16), we also have
$\lambda_{j}=b_{i1}+b_{i2}\theta_{j}^{-1}+\cdots+b_{ik}\theta_{j}^{-(k-1)}~~(j=1,2,\cdots,k).$ (17)
Thus, $\lambda_{2},\cdots,\lambda_{k}$ are the conjugates of $\lambda_{1}$ corresponding to $\theta_{2},\cdots,\theta_{k}$. Since $\det B=\pm 1$ and (15), we have $\lambda_{1}\lambda_{2}\cdots\lambda_{k}=\pm 1$, which implies that the norm of $\lambda_{1}$ is $\pm 1$, and hence, $\lambda_{1}$ is a unit.
Conversely, if $\lambda_{1}$ is a unit satisfying (15). Define $B$ with the cojugates $\lambda_{2},\cdots,\lambda_{k}$ of $\lambda_{1}$ by (16). Then, $B$ is a integer matrix since (15), (17) and the representations as (17) are unique. Moreover it commutes with $C_{P}$ and $\det B=\pm 1$, we have $B\in Saut_{P}$. Thus, we have a bijection bewtween $Saut_{P}$ and the set of units in $\mathbb{Q}(\theta_{1})$ satisfying (15). It is clear that this is an isomorphism with respect to the multiplications.
(ii) By the unit theorem of Dirichlet (Boreivich and Shafarevich [1]), the set of units in $\mathbb{Q}(\theta_{1})$ is a direct product of $c_{1}+c_{2}$ number of cyclic groups of infinite orders with a finite group, where $k_{1},2k_{2}$ are the numbers of real roots and complex roots of $P(z)=0$, respectively. Since the set of units satisfying (15) is a subpgroup of it which is isomorphic to $Saut_{P}$, $Saut_{P}$ is a direct product of $l$ number of cyclic groups of infinite orders with a finite group, where $l\leq k_{1}+k_{2}-1$. ∎
Example 3.
Let $P(z)=2z^{2}-2z-3$. It has the roots $(1\pm\sqrt{7})/2$. Let $\theta_{1}=(1+\sqrt{7})/2$. The fundamental units in $\mathbb{Q}(\sqrt{7})$ are $8\pm 3\sqrt{7}$. Let $\alpha=8+3\sqrt{7}$. Then, it satisfies (15) since
$\alpha=6\theta_{1}+5=9\theta_{1}^{-1}+11.$
Let $B\in Saut_{P}$ be the matrix corresponding to $\lambda_{1}=\alpha$ in (15). Then we have $B=\begin{pmatrix}11&9\\ 6&5\end{pmatrix}$. Hence, $Saut_{P}/\pm I_{2}=\left\{\begin{pmatrix}11&9\\ 6&5\end{pmatrix}^{n};~n\in\mathbb{Z}\right\}$ follows.
Example 4.
Let $P(z)=z^{2}-125$. It has the roots $\pm 5\sqrt{5}$ and the fundamental units in $\mathbb{Q}(\sqrt{5})$ are $\frac{1\pm\sqrt{5}}{2}$. Let $\alpha=\frac{1+\sqrt{5}}{2}$. The smallest power $k$ such that $\alpha^{k}$ satisfies (15) is 8. Let $\lambda_{1}:=\alpha^{8}=682+305\sqrt{5}$. Let $B\in Saut_{P}$ be the matrix corresponding to $\lambda_{1}$ in (15). Then we have $B=\begin{pmatrix}682&7625\\ 61&682\end{pmatrix}$. Hence,
$Saut_{P}/\pm I_{2}=\left\{\begin{pmatrix}682&7625\\ 61&682\end{pmatrix}^{n};n\in\mathbb{Z}\right\}.$
Example 5.
Let $\theta_{1},\theta_{2},\theta_{3}$ be the roots of $P(z)=2z^{3}-3z^{2}-3z+3$ which are $3$ distinct real numbers. Then, we have
$C_{P}=\begin{pmatrix}0&1&0\\ 0&0&1\\ -2/3&1&1\end{pmatrix}$.
We have $\mathbb{Z}[3\alpha]\subsetneq\mathbb{O}_{K}=\mathbb{Z}+3\alpha\mathbb{Z}+3\alpha^{2}\mathbb{Z}$ and $\mathbb{O}_{K}^{*}=\{\pm\epsilon_{1}^{n}\epsilon_{2}^{m}\ |\ m,n\in\mathbb{Z}\}$ with
$\epsilon_{1}=1-3\alpha^{2},\ \epsilon_{2}=5-6\alpha^{2}$
where $\epsilon_{1},\epsilon_{2}\in\mathbb{O}_{K}^{*}$ and $\epsilon_{1},\epsilon_{2}\not\in\mathbb{Z}[3\alpha]$. In this case, we have
$B_{1}=\iota(\epsilon_{1})=E_{3}-3C_{P}^{2}=\begin{pmatrix}1&0&-3\\ 2&-2&-3\\ 2&-1&-5\\ \end{pmatrix}$,
$B_{2}=\iota(\epsilon_{2})=5E_{3}-6C_{P}^{2}=\begin{pmatrix}5&0&-6\\ 4&-1&-6\\ 4&-2&-7\\ \end{pmatrix}$
and both of them are in $M_{3\times 3}(\mathbb{Z})$. Therefore
$Saut=\{\pm B_{1}^{n}B_{2}^{m}\ |\ m,n\in\mathbb{Z}\}.$
Considering the inverse $1/\alpha$, from regular representation by $\{1,1/\alpha,1/\alpha^{2}\}$, we have
$\tilde{C}_{P}=\begin{pmatrix}0&1&0\\ 0&0&1\\ -3/2&3/2&3/2\\ \end{pmatrix}$.
We have $\mathbb{Z}[2/\alpha]\subsetneq\mathbb{O}_{K}=\mathbb{Z}+2/\alpha\mathbb{Z}+(1/\alpha+2/\alpha^{2})\mathbb{Z}$ and
$\epsilon_{1}=-5-1/\alpha+2/\alpha^{2},\ \epsilon_{2}=-7-2/\alpha+4/\alpha^{2}$
and notice that $\epsilon_{2}\in\mathbb{Z}[2/\alpha]$. Thus
$\tilde{B}_{1}=\iota(\epsilon_{1})=-5E_{3}-\tilde{C}_{P}+2\tilde{C}_{P}^{2}=\begin{pmatrix}-5&-1&2\\ -3&-2&2\\ -3&0&1\\ \end{pmatrix}$,
$\tilde{B}_{2}=\iota(\epsilon_{2})=-7E_{3}-2\tilde{C}_{P}+4\tilde{C}_{P}^{2}=\begin{pmatrix}-7&-2&4\\ -6&-1&4\\ -6&0&5\\ \end{pmatrix}$
and both of them are in $M_{3\times 3}(\mathbb{Z})$. Therefore
$\widetilde{Saut}=\{\pm\tilde{B}_{1}^{n}\tilde{B}_{2}^{m}\ |\ m,n\in\mathbb{Z}\}.$
Example 6.
Let $\alpha$ by a root of $2x^{3}-3x^{2}+2x+3$ with $r_{1}=1$ and $r_{2}=1$. Similarly we have
$C_{P}=\begin{pmatrix}0&1&0\\ 0&0&1\\ -3/2&-1&3/2\\ \end{pmatrix}$.
We have $\mathbb{Z}[2\alpha]\subsetneq\mathbb{O}_{K}=\mathbb{Z}+2\alpha\mathbb{Z}+(1/2+3\alpha/2+\alpha^{2})\mathbb{Z}$ and $\mathbb{O}_{K}^{*}=\{\pm\epsilon^{n}\ |\ n\in\mathbb{Z}\}$ with
$\epsilon=\frac{1}{2}+\frac{1}{2}\alpha-\alpha^{2}$ (18)
and $\epsilon\not\in\mathbb{Z}[2\alpha]$. Let
$V=\iota(\epsilon)=\frac{1}{2}E_{3}+\frac{1}{2}C_{P}-C_{P}^{2}=\begin{pmatrix}1/2&1/2&-1\\ 3/2&3/2&-1\\ 3/2&5/2&0\\ \end{pmatrix}$.
Then we see that
$B=V^{3}=\begin{pmatrix}-4&-5&2\\ -3&-6&-2\\ 3&-1&-9\\ \end{pmatrix}$
belongs to $M_{3\times 3}(\mathbb{Z})$ and we have
$Saut=\{\pm B^{n}\ |\ n\in\mathbb{Z}\}.$
We see $\epsilon^{3}=-4-5\alpha+2\alpha^{2}\in\mathbb{Z}[\alpha]\setminus\mathbb{Z}[2\alpha]$. We claim that $\epsilon$ and $\epsilon^{2}$ do not belongs to $\mathbb{Z}[\alpha]$. Assume that $\epsilon\in\mathbb{Z}[\alpha]$ and take an expression $\epsilon=\sum_{i=0}^{\ell}x_{i}\alpha^{i}$ with $x_{i}\in\mathbb{Z}$ with the minimum $\ell$. From (18), we know that $\ell\geq 3$. Then $\rm{Tr}(C_{P}^{n})\ (n=0,1,\dots)$ form a sequence
[…]
which satisfies the linear recurrence
$a_{n+3}=\frac{3}{2}a_{n+2}-a_{n+1}-\frac{3}{2}a_{n}$
and we can show that its denominator is $2^{n}\ (n=0,1,\dots)$. Since
$2=\rm{Tr}V=\sum_{i=0}^{\ell}x_{i}\rm{Tr}(C_{P}^{i})$
we know that $x_{\ell}$ is even. However then using
$\alpha^{\ell}=\frac{3}{2}\alpha^{\ell-1}-\alpha^{\ell-2}-\frac{3}{2}\alpha^{\ell-3}$
we have an expression of $\epsilon$ in $\mathbb{Z}[\alpha]$ of length $\ell-1$, which is a contradiction. Similarly we see $\epsilon^{2}\not\in\mathbb{Z}[\alpha]$. The claim is shown.
Example 7.
Let $\alpha$ be a root of $3x^{4}-2x^{3}+x^{2}+x-2$ with $r_{1}=2$ and $r_{2}=1$. Then
$C_{P}=\begin{pmatrix}0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ 2/3&-1/3&-1/3&2/3\\ \end{pmatrix}$.
We have $\mathbb{O}_{K}=\mathbb{Z}+3\alpha\mathbb{Z}+(\alpha+3\alpha^{2})\mathbb{Z}+(2\alpha+\alpha^{2}+3\alpha^{3})\mathbb{Z}$ and
$\mathbb{O}_{K}^{*}=\{\pm\epsilon_{1}^{m}\epsilon_{2}^{n}\ |\ m,n\in\mathbb{Z}\}$
with
$\epsilon_{1}=1-2\alpha-2\alpha^{2}+3\alpha^{3},\ \epsilon_{2}=361+237\alpha+93\alpha^{2}+468\alpha^{3}.$
In this case, we have
$B_{1}=\iota(\epsilon_{1})=\begin{pmatrix}1&-2&-2&3\\ 2&0&-3&0\\ 0&2&0&-3\\ -2&1&3&-2\\ \end{pmatrix}$,
$B_{2}=\iota(\epsilon_{2})=\begin{pmatrix}361&237&93&468\\ 312&205&81&405\\ 270&177&70&351\\ 234&153&60&304\\ \end{pmatrix}$
and
$Saut=\{\pm B_{1}^{n}B_{2}^{m}\ |\ m,n\in\mathbb{Z}\}.$
7 $Saut_{P}$ in the case $\deg P=2,3$
Theorem 8.
[…]
(i) Assume $D>0$ and let the roots of $P(z)=0$ be $\theta_{1},\theta_{2}$. Then the following conditions for an integer matrix $B=\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)$ are equivalent:
(i-1) $B\in Saut_{P}$,
[…]
(i-3)  $B=\left(\begin{array}[]{cc}c_{0}&a_{0}c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right)$ for some $c_{0},c_{1}\in\mathbb{Z}$ such that
$(2c_{0}-a_{1}c_{1})^{2}-Dc_{1}^{2}=\pm 4,$ (19)
(i-4) $B$ satisfies that $pq^{\prime}-p^{\prime}q=\pm 1$ and
$\frac{p\theta_{i}+p^{\prime}}{q\theta_{i}+q^{\prime}}=\theta_{i}~~(i=1,2).$ (20)
Moreover, $Saut_{P}/\{\pm I_{2}\}$ is an infinite cyclic group if and only if $D$ is a non-square integer. In addition, if $D\in\{1,4\}$, then it is a cyclic group of order $2$. Otherwise, it is the trivial group.
(ii) If $D=0$, then $Saut_{P}/\{\pm I_{2}\}$ is an infinite cyclic group.
(iii) Assume $D<0$. Then $Saut_{P}/\{\pm I_{2}\}$ is the trivial group except for the case $D\in\{-3,-4\}$. If D=3D=-3, then $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order $3$. If D=4D=-4, then $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order $2$.
Remark 3.
The equation (19) is known as the Pell’s equation, i.e., Diophantine equation of the form
$w^{2}-Dv^{2}=\pm 4,\ w,v\in Z,~\mbox{for a given integer}\,D.$ (21)
The main tool to solve it is the theory of continued fraction (see [2, 7], for example).
Proof.
(i) Assume $D>0$.
By Theorem 7, we already know that (i-1) and (i-2) are equivalent.
Assume (i-2). Then we have
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}\theta_{1}^{2}&\theta_{2}^{2}\\ \theta_{1}&\theta_{2}\end{array}\right)\left(\begin{array}[]{cc}\lambda_{1}&0\\ 0&\lambda_{2}\end{array}\right)\left(\begin{array}[]{cc}\theta_{1}^{2}&\theta_{2}^{2}\\ \theta_{1}&\theta_{2}\end{array}\right)^{-1}$ (22)
with some $\lambda_{i}~(i=1,2)$ such that $\lambda_{1}\lambda_{2}=\pm 1$. It follows from (22) that $\lambda_{i}=c_{0}+c^{\prime}_{1}\theta_{i}^{-1}$ with $c_{0}=p,~c^{\prime}_{1}=p^{\prime}$. Combining with $\theta_{1}\theta_{2}=\frac{a_{0}}{a_{2}},~\theta_{1}+\theta_{2}=-\frac{a_{1}}{a_{2}}$, we thus have
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}c_{0}&c^{\prime}_{1}\\ -\frac{a_{2}}{a_{0}}c^{\prime}_{1}&c_{0}-\frac{a_{1}}{a_{0}}c^{\prime}_{1}\end{array}\right).$
As $B$ is a integer matrix, $\frac{a_{2}}{a_{0}}c^{\prime}_{1}$ and $\frac{a_{1}}{a_{0}}c^{\prime}_{1}$ are integers. Since $a_{2}z^{2}+a_{1}z+a_{0}$ is a primitive polynomial, this is possible only if $c^{\prime}_{1}$ is a multiple of $a_{0}$. Putting $c_{1}=\frac{c^{\prime}_{1}}{a_{0}}\in\mathbb{Z}$, we have $c_{1}^{\prime}=a_{0}c_{1}$ and
$B=\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}c_{0}&a_{0}c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right).$
Clearly, $pq^{\prime}-p^{\prime}q=\pm 1$ is equivalent to (19). Thus, (i-2) implies (i-3).
Conversely, if $B=\left(\begin{array}[]{cc}c_{0}&a_{0}c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right)$, then $BC_{P}=C_{P}B$ follows. Thus, (i-3) implies (i-2).
Let $B\in Saut_{p}$. Then by (16), we have
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)\left(\begin{array}[]{cc}\theta_{1}^{2}&\theta_{2}^{2}\\ \theta_{1}&\theta_{2}\end{array}\right)=\left(\begin{array}[]{cc}\theta_{1}^{2}&\theta_{2}^{2}\\ \theta_{1}&\theta_{2}\end{array}\right)\left(\begin{array}[]{cc}\lambda_{1}&0\\ 0&\lambda_{2}\end{array}\right).$
for some $\lambda_{i}~(i=1,2)$ with $\lambda_{1}\lambda_{2}=\pm 1$. It follows that
$\frac{p\theta_{i}+p^{\prime}}{q\theta_{i}+q^{\prime}}=\theta_{i}\mbox{ and }pq^{\prime}-p^{\prime}q=\pm 1~~(i=1,2).$
Thus, (i-1) implies (i-4).
Assume (i-4). Then, therer exists $\lambda_{1},\lambda_{2}$ such that
$\left(\begin{array}[]{cc}p&p^{\prime}\\ q&q^{\prime}\end{array}\right)\left(\begin{array}[]{c}\theta_{i}^{2}\\ \theta_{i}\end{array}\right)=\lambda_{i}\left(\begin{array}[]{c}\theta_{i}^{2}\\ \theta_{i}\end{array}\right)~~(i=1,2).$
Hence, we have (16) and $\lambda_{1}\lambda_{2}=pq^{\prime}-p^{\prime}q=\pm 1$. Thus, (i-4) implies (i-2) and the first half of statement (i) is proved.
Let $D>0$. Consider the equation (19), it always has trivial solutions $c_{0}=\pm 1,~c_{1}=0$, that correspond to $B=\pm I_{2}$. If $D$ is a square integer, then it is easy to see that (19) has a nontrivial solution only if $D=4$, since the difference of 2 square integers is 4 only if they are 0 and 4. Assume that $Saut_{P}/\{\pm I_{2}\}$ has an element, say $B$ of order $2$. This implies that
$B$ $=\left(\begin{array}[]{cc}\theta_{1}^{2}&\theta_{2}^{2}\\ \theta_{1}&\theta_{2}\end{array}\right)\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)\left(\begin{array}[]{cc}\theta_{1}^{2}&\theta_{2}^{2}\\ \theta_{1}&\theta_{2}\end{array}\right)^{-1}$
$=\frac{1}{\theta_{1}-\theta_{2}}\left(\begin{array}[]{cc}\theta_{1}+\theta_{2}&-2\theta_{1}\theta_{2}\\ 2&-(\theta_{1}+\theta_{2})\end{array}\right)=\frac{1}{\theta_{1}-\theta_{2}}\left(\begin{array}[]{cc}-\frac{a_{1}}{a_{2}}&-\frac{2a_{0}}{a_{2}}\\ 2&\frac{a_{1}}{a_{2}}\end{array}\right)$
is an integer matrix. Therefore $\theta_{1}-\theta_{2}$ is rational, and hence, $D$ is a square integer.
Thus, if $D$ is a nonsquare integer, then $Saut_{P}/\{\pm I_{2}\}$ except for the unit contains only elements of infinite order, and later we prove that it does contain elements of infinite order. Let $\eta$ be the mapping defined on $Saut_{P}$ such that for $B\in Saut_{P}$, $\eta(B)=(\lambda_{1},\lambda_{2})$, where $(\lambda_{1},\lambda_{2})$ satisfies (22). Then $\eta(B)\in(\mathbb{R}\setminus\{0\})^{2}$ and $Image(\eta)$ is a discrete multiplicative subgroup of $(\mathbb{R}\setminus\{0\})^{2}$. Let $\tilde{\lambda}$ be the smallest element in $\{\lambda_{1}\in(1,\infty);~(\lambda_{1},\lambda_{2})\in Image(\eta)\}$. Since $(1,-1)\notin\eta(B)$, just one of $(\tilde{\lambda},\tilde{\lambda}^{-1})$ or $(\tilde{\lambda},-\tilde{\lambda}^{-1})$ is in $Image(\eta)$. This element generates $\tilde{\lambda}$.
Let $D>0$ be a nonsquare integer. We prove that $Saut_{P}/\{\pm I_{2}\}$ contains nontrivial elements using the continued fraction expansion of $\theta_{1}$, which is eventually periodic. Let it be
$\theta_{1}=[c_{0};c_{1},\cdots,c_{k-1},(e_{0},e_{1}\cdots,e_{m-1})^{\infty}].$
Then,
$\frac{C}{G}=[c_{0};c_{1},\cdots,c_{k-1}],~\frac{C^{\prime}}{G^{\prime}}=[c_{0};c_{1},\cdots,c_{k-2}]$
$\frac{E}{F}=[e_{0};e_{1},\cdots,e_{m-1}],~\frac{E^{\prime}}{F^{\prime}}=[e_{0};e_{1},\cdots,e_{m-2}],$
are irreducible fractions with nonnegative denominators, where if $k=0$, that is, the continued fraction expansion is purely periodic, then we put
$\left(\begin{array}[]{cc}C&C^{\prime}\\ G&G^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)$
and if $k=1$ or $m=1$, we put
$\left(\begin{array}[]{cc}C&C^{\prime}\\ G&G^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}c_{0}&1\\ 1&0\end{array}\right)~\mbox{and}~\left(\begin{array}[]{cc}E&E^{\prime}\\ F&F^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}e_{0}&1\\ 1&0\end{array}\right).$
Finally, define $2\times 2$-matrix $\left(\begin{array}[]{cc}p_{n}&p^{\prime}_{n}\\ q_{n}&q^{\prime}_{n}\end{array}\right)~(n\in\mathbb{Z})$ by
$\left(\begin{array}[]{cc}p_{n}&p^{\prime}_{n}\\ q_{n}&q^{\prime}_{n}\end{array}\right)=\left(\begin{array}[]{cc}C&C^{\prime}\\ G&G^{\prime}\end{array}\right)\left(\begin{array}[]{cc}E&E^{\prime}\\ F&F^{\prime}\end{array}\right)^{n}\left(\begin{array}[]{cc}C&C^{\prime}\\ G&G^{\prime}\end{array}\right)^{-1}.$
Then, they are different each other and the totality of elements in $Saut_{P}/\{\pm I_{2}\}$. Hence, $Saut_{P}/\{\pm I_{2}\}$ is an infinite cyclic group.
Let $D>0$ be a square. By (i-3), for an integer matrix $B=\left(\begin{array}[]{cc}c_{0}&a_{0}c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right)$, $B\in Saut_{P}$ if and only if for some $c_{0},c_{1}\in\mathbb{Z}$ such that (19) holds, That is,
$(2c_{0}-a_{1}c_{1})^{2}-Dc_{1}^{2}=\pm 4.$
It always has solution $c_{0}=\pm 1,~c_{1}=0$ corresponding to $B=\pm I$. The other solutions come from the case $c_{1}\neq 0$. Since the difference between 2 square numbers becomes 4 only if one is 0 and the other is 4, the solution of (19) with $c_{1}\neq 0$ is limited to the following 2 cases:
$\left\{\begin{array}[]{c}D=1\\ c_{1}=\pm 2\\ a_{1}=\pm c_{0}\end{array}\right.(\pm~\mbox{correspondinfly}),~\mbox{and}~\left\{\begin{array}[]{c}D=4\\ c_{1}=\pm 1\\ a_{1}=\pm 2c_{0}\end{array}\right.(\pm~\mbox{correspondinfly})$
In the first case, we have $B=\pm\left(\begin{array}[]{cc}a_{1}&2a_{0}\\ -2a_{2}&-a_{1}\end{array}\right)$ which satisfies that $B^{2}=I_{2}$ in vertue of $a_{1}^{2}-4a_{0}a_{2}=1$. Thus, $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order 2.
In the second case, note that $a_{1}$ is even and we have $B=\pm\left(\begin{array}[]{cc}\frac{a_{1}}{2}&a_{0}\\ -a_{2}&-\frac{a_{1}}{2}\end{array}\right)$ which satisfies that $B^{2}=I_{2}$ in vertue of $a_{1}^{2}-4a_{0}a_{2}=4$. Thus, Thus, $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order 2.
(ii) Let $D=0$. Then by (10), $B\in Saut_{P}$ if and only if
$B=\left(\begin{array}[]{cc}p&p^{\prime}\\ -a_{2}p^{\prime}&p-a_{1}p^{\prime}\end{array}\right)\mbox{ with }\det B=\left(p-\frac{a_{1}p^{\prime}}{2}\right)^{2}=1.$
Since $D=a_{1}^{2}-4a_{2}=0$, $a_{1}$ is even. Put $a_{1}=2c$. Then, $p-cp^{\prime}=\pm 1$ and
$B=\left(\begin{array}[]{cc}\pm 1+cp^{\prime}&p^{\prime}\\ -c^{2}p^{\prime}&\pm 1-cp^{\prime}\end{array}\right)\in Saut_{P}~(\forall p^{\prime}\in\mathbb{Z})~(\mbox{$\pm$ correspondingly}).$
Thus, we have Case 3 and $Saut_{P}/\{\pm I_{2}\}$ is an infinite cyclic group.
(iii) Note that $D\equiv 1~\mbox{or}~0~(\rm{mod}~4)$. Hence, if $D<0$, then $D\leq-3$. Let $B\in Saut_{P}$. By (i-3), $B=\left(\begin{array}[]{cc}c_{0}&a_{0}c_{1}\\ -a_{2}c_{1}&c_{0}-a_{1}c_{1}\end{array}\right)$ for some $c_{0},c_{1}\in\mathbb{Z}$ with (19). Therefore, $(2c_{0}-a_{1}c_{1})^{2}-Dc_{1}^{2}=4$. This always has a trivial solution that $c_{1}=0,~c_{0}=\pm 1$, corresponding to $B=\pm I_{2}$. We look for the other solutions.
There are 2 cases where the nontrivial sdolutions exist.
(Case 1) D=3D=-3 and $\left\{\begin{array}[]{c}2c_{0}-a_{1}c_{1}=\pm 1\\ c_{1}=\pm 1\end{array}\right.~\mbox{($\pm$ independently)}$.
In this case, $a_{1}$ is an odd number.
If $c_{1}=1$, then $B\in Saut_{P}$ is one of the following:
$B_{1}=\left(\begin{array}[]{cc}\frac{a_{1}+1}{2}&a_{0}\\ -a_{2}&\frac{-a_{1}+1}{2}\end{array}\right)~\mbox{or}~B_{2}=\left(\begin{array}[]{cc}\frac{a_{1}-1}{2}&a_{0}\\ -a_{2}&\frac{-a_{1}-1}{2}\end{array}\right).$
If $c_{1}=-1$, then $B\in Saut_{P}$ is one of the following:
$B_{3}=\left(\begin{array}[]{cc}\frac{-a_{1}-1}{2}&-a_{0}\\ a_{2}&\frac{a_{1}+1}{2}\end{array}\right)~\mbox{or}~B_{4}=\left(\begin{array}[]{cc}\frac{-a_{1}+1}{2}&-a_{0}\\ a_{2}&\frac{a_{1}-1}{2}\end{array}\right).$
It is easy to check using $a_{1}^{2}-4a_{0}a_{2}=-3$ that
$B_{1}^{2}=B_{2},~B_{1}^{3}=-I_{2}~\mbox{and}~B_{3}=-B_{1},~B_{4}=-B_{2}.$
This implies that $Saut_{P}/\{\pm I_{2}\}$ is cyclic group of order 3.
(Case 2) D=4D=-4. In this case, just same as the case $D=4$, we have $Saut_{P}/\pm 1$ is a cyclic group of order 2. ∎
We now give several examples to illustrate Theorem 8.
Example 8.
Let $P(z)=5z^{2}-2$ and $\theta=\frac{\sqrt{10}}{5}$ be one of the roots of $P(z)=0$. It has the continued fraction expansion
$\theta=[0;1,\underbrace{1,1,2},\underbrace{1,1,2},\underbrace{1,1,2},\cdots].$
Then let
$\left(\begin{array}[]{cc}C&C^{\prime}\\ G&G^{\prime}\end{array}\right)=\left(\begin{array}[]{cc}1&0\\ 1&1\end{array}\right)$
$\frac{E}{F}:=[1;1,2]=1+\frac{1}{1+\frac{1}{2}}=\frac{5}{3}~,~~\frac{E^{\prime}}{F^{\prime}}:=[1;1]=1+\frac{1}{1}=\frac{2}{1}.$
Then, we have
$Saut_{P}/\{\pm I_{2}\}=\left\{\left(\begin{array}[]{cc}1&0\\ 1&1\end{array}\right)\left(\begin{array}[]{cc}5&2\\ 3&1\end{array}\right)^{n}\left(\begin{array}[]{cc}1&0\\ 1&1\end{array}\right)^{-1};~n\in\mathbb{Z}\right\}.$
If $n=1$ in the above, we have $\left(\begin{array}[]{cc}3&2\\ 5&3\end{array}\right)$. Applying it to
$(\cdots;\frac{3}{7},\frac{5}{7},\frac{4}{7},\frac{2}{7},\frac{3}{7},\frac{5}{7},\cdots)\in\Omega_{P},$
we get
$(\cdots;\frac{5}{7},\frac{2}{7},\frac{2}{7},\frac{5}{7},\frac{5}{7},\frac{2}{7},\cdots)\in\Omega_{P}.$
Example 9.
Let $P(z)=z^{2}-2z+1$ with $D=0$. Then by Theorem 8, $Saut_{P}=\{\pm B^{n};n\in\mathbb{Z}\}$ with $B=\left(\begin{array}[]{cc}0&1\\ -1&2\end{array}\right)$. Applying it to $(x_{n};n\in\mathbb{Z})\in\Omega_{P}$, we get $(y_{n};n\in\mathbb{Z})\in\Omega_{P}$ with
$\left(\begin{array}[]{c}y_{n}\\ y_{n+1}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ -1&2\end{array}\right)\left(\begin{array}[]{c}x_{n}\\ x_{n+1}\end{array}\right)=\left(\begin{array}[]{c}x_{n+1}\\ -x_{n}+2x_{n+1}\end{array}\right).$
Since $y_{n+1}=-x_{n}+2x_{n+1},~y_{n+2}=-y_{n}+2y_{n+1}=-2x_{n}+3x_{n+1}$, we have
$\left(\begin{array}[]{c}y_{n}\\ y_{n+1}\end{array}\right)=\left(\begin{array}[]{cc}0&1\\ -1&2\end{array}\right)\left(\begin{array}[]{c}x_{n+1}\\ -x_{n}+2x_{n+1}\end{array}\right)=\left(\begin{array}[]{c}-x_{n}+2x_{n+1}\\ -2x_{n}+3x_{n+1}\end{array}\right)$
as expected.
Example 10.
Let $P(z)=z^{2}+1\in\mathbb{Z}[z]$ with D=4D=-4. Then it is easy to show that $Saut_{P}/\{\pm I_{2}\}=\left\{I_{2},\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)\right\}$. Since $\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)^{2}=I_{2}$, $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order 2.
We now discuss the case of degree 3. It’s much harder than the case of degree 2 and we have a result only for a special case.
Theorem 9.
(i) Assume that the roots $\theta_{1},\theta_{2},\theta_{3}$ of
$P(z)=a^{\prime}_{3}z^{3}+a^{\prime}_{2}z^{2}+a^{\prime}_{1}z+a^{\prime}_{0}\in\mathbb{Z}[z],~a_{0},a_{3}\neq 0$
are simple. We put $a_{3}=\frac{a^{\prime}_{3}}{a^{\prime}_{0}},~a_{2}=\frac{a^{\prime}_{2}}{a^{\prime}_{0}},~a_{1}=\frac{a^{\prime}_{1}}{a^{\prime}_{0}}$. Then $B\in Saut_{P}$ if and only if
$B=\left(\begin{array}[]{ccc}c_{0}&c_{1}&c_{2}\\ -a_{3}c_{2}&c_{0}-a_{2}c_{2}&c_{1}-a_{1}c_{2}\\ -a_{3}c_{1}+a_{1}a_{3}c_{2}&-a_{2}c_{1}+(a_{1}a_{2}-a_{3})c_{2}&c_{0}-a_{1}c_{1}+(a_{1}^{2}-a_{2})c_{2}\end{array}\right)$
for some $c_{0},c_{1},c_{2}\in\mathbb{Z}$ such that
$\begin{split}~~~\det B&=c_{0}^{3}-a_{3}c_{1}^{3}+a_{3}^{2}c_{2}^{3}-a_{1}c_{0}^{2}c_{1}+(a_{1}^{2}-4a_{2})c_{0}^{2}c_{2}+a_{2}c_{0}c_{1}^{2}\\ &+(a_{2}^{2}-2a_{1}a_{3})c_{0}c_{2}^{2}+a_{1}a_{3}c_{1}^{2}c_{2}-a_{2}a_{3}c_{1}c_{2}^{2}+(3a_{3}-a_{1}a_{2})c_{0}c_{1}c_{2}\\ &=\pm 1.\end{split}$
(ii) Let $P(z)=-az^{3}+1\in\mathbb{Z}[z]$. Then, $B\in Saut_{P}$ if and only if $B=\left(\begin{array}[]{ccc}c_{0}&c_{1}&c_{2}\\ ac_{2}&c_{0}&c_{1}\\ ac_{1}&ac_{2}&c_{0}\end{array}\right)$ for some $c_{0},c_{1},c_{2}\in\mathbb{Z}$ such that
$\det B=c_{0}^{3}+ac_{1}^{3}+a^{2}c_{2}^{3}-3ac_{0}c_{1}c_{2}=\pm 1.$ (23)
(iii) Let $a=1$ in (i). Then, $Saut_{P}/\{\pm I_{3}\}$ is a cyclic group of order $3$ spanned by $\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)$.
Proof.
(i) By hypothesis, $\theta_{i}^{-1}(i=1,2,3)$ is the roots of the reciprocal polynomial of $p(z)$, say $h(z)$, i.e., $h(z)=z^{3}+a_{1}z^{2}+a_{2}z+a_{3}$. Put $\tau_{1}:=\frac{1}{\theta_{1}}+\frac{1}{\theta_{2}}+\frac{1}{\theta_{3}}$, $\tau_{2}:=\frac{1}{\theta_{1}\theta_{2}}+\frac{1}{\theta_{1}\theta_{3}}+\frac{1}{\theta_{2}\theta_{3}}$, $\tau_{3}:=\frac{1}{\theta_{1}\theta_{2}\theta_{3}}$.
Applying theorem 5 with $k=3$,
$B=\left(\begin{array}[]{ccc}\theta_{1}^{3}&\theta_{2}^{3}&\theta_{3}^{3}\\ \theta_{1}^{2}&\theta_{2}^{2}&\theta_{3}^{2}\\ \theta_{1}&\theta_{2}&\theta_{3}\end{array}\right)\left(\begin{array}[]{ccc}\lambda_{1}&0&0\\ 0&\lambda_{2}&0\\ 0&0&\lambda_{3}\end{array}\right)\left(\begin{array}[]{ccc}\theta_{1}^{3}&\theta_{2}^{3}&\theta_{3}^{3}\\ \theta_{1}^{2}&\theta_{2}^{2}&\theta_{3}^{2}\\ \theta_{1}&\theta_{2}&\theta_{3}\end{array}\right)^{-1}$
where $\lambda_{i}=c_{0}+c_{1}\theta_{i}^{-1}+c_{2}\theta_{i}^{-2}~(i=1,2,3).$ A careful calculation from above gives
$B=\left(\begin{array}[]{ccc}c_{1}&c_{2}&c_{3}\\ c_{2}\tau_{3}&c_{0}-c_{2}\tau_{2}&c_{1}+c_{2}\tau_{1}\\ c_{1}\tau_{3}+c_{2}\tau_{1}\tau_{3}&-c_{1}\tau_{2}-c_{2}\tau_{1}\tau_{2}+c_{2}\tau_{3}&c_{0}+c_{1}\tau_{1}+c_{2}(\tau_{1}^{2}-\tau_{2})\end{array}\right)$
From Vieta’s formulas, we have $\tau_{1}=-a_{1},\tau_{2}=a_{2},\tau_{3}=-a_{3}.$ The second statement follows just by calculating $\det B$, so (i) follows.
(ii) From $BC_{P}=C_{P}B$, the expression of B in (i) follows. Or applying (i) with $a_{1}=a_{2}=0,a_{3}=-a.$
(iii) By (ii), $B\in Saut_{P}$ if and only if
$\det B=c_{0}^{3}+c_{1}^{3}+c_{2}^{2}-3c_{0}c_{1}c_{2}$
$=(c_{0}+c_{1}+c_{2})(c_{0}^{2}+c_{1}^{2}+c_{2}^{2}-c_{0}c_{1}-c_{1}c_{2}-c_{2}c_{1})$
$=\frac{1}{2}(c_{0}+c_{1}+c_{2})((c_{0}-c_{1})^{2}+(c_{1}-c_{2})^{2}+(c_{2}-c_{0})^{2})$
Therefore, by the second line above $\lambda_{1}\lambda_{2}\lambda_{3}=\pm 1$ holds only if $c_{0}+c_{1}+c_{2}=\pm 1$. Hence, by the third line above $(c_{0}-c_{1})^{2}+(c_{1}-c_{2})^{2}+(c_{2}-c_{0})^{2}=2$. This is possible only if two of the terms in the left hand side is 1 and the other is 0. Without loss of generality, assume that $c_{0}=c_{1}$ and $c_{2}=c_{0}\pm 1$. Then, we have $c_{0}+c_{1}+c_{2}=3c_{0}\pm 1$. Since $c_{0}+c_{1}+c_{2}=\pm 1$, we must have $c_{0}=0$, and hence, $(c_{0},c_{1},c_{2})=(0,0,\pm 1)$. Thus, $\lambda_{1}\lambda_{2}\lambda_{3}=\pm 1$ is equivalent to $(c_{0},c_{1},c_{2})\in\{(0,0,\pm 1),(0,\pm 1,0),(\pm 1,0,0)\}$, and hence, $(\lambda_{1},\lambda_{2},\lambda_{3})\in\{(\pm 1,\pm 1,\pm 1),(\pm\omega,\pm\omega,\pm\omega),(\pm\omega^{2},\pm\omega^{2},\pm\omega^{2})\}$. Thus, $Saut_{P}/\{\pm I_{2}\}$ is a cyclic group of order 3. ∎
Example 11.
Let $P(z)=-z^{3}+1\in\mathbb{Z}[z]$ and $\omega=\frac{-1+\sqrt{3}}{2}$. Then, by Theorem 7, $B=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)$. Since we can easily verify that
$B^{2}=\left(\begin{array}[]{ccc}0&0&1\\ 1&0&0\\ 0&1&0\end{array}\right)\neq I_{3},~B^{3}=I_{3}.$
Therefore, $Saut_{P}/\{\pm I_{3}\}$ is a cyclic group of order $3$ and it is spanned by $B=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)$. Furthermore Let $X=(x_{n};n\in\mathbb{Z})\in\Omega_{P}$ and $\varphi_{B}(X)=(y_{n};n\in\mathbb{Z})$. Then,
$\left(\begin{array}[]{c}y_{0}\\ y_{1}\\ y_{2}\end{array}\right)=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)\left(\begin{array}[]{c}x_{0}\\ x_{1}\\ x_{2}\end{array}\right)=\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ x_{0}\end{array}\right).$
Since $x_{3}=x_{0},~y_{3}=y_{0}$, we have
$\left(\begin{array}[]{c}y_{1}\\ y_{2}\\ y_{3}\end{array}\right)=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right)\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ x_{3}\end{array}\right)=\left(\begin{array}[]{c}x_{2}\\ x_{3}\\ x_{1}\end{array}\right)=\left(\begin{array}[]{c}x_{2}\\ x_{0}\\ x_{1}\end{array}\right)$
as expected.
8 Streams over Galois fields
In the last section, we investigate the streams over the Galois field $GF(q^{k})=GF(q)(\theta)$, where $\theta$ is a root of an irreducible polynomial $P(z)\in GF(q)[z]$ of degree $k$. The zeros $\Omega_{P}^{q}$ of $ST(GF(q^{k}))$ by the action of $P(Z)$ is discussed. Note that any element in $\Omega_{P}^{q}$ is cyclic with period $q^{k}-1$.
Theorem 10.
Let $P(z)\in GF(q)[z]$ be an irreducible polynomial of degree $k\geq 1$. Let $\Omega_{P}^{q}$ be the zeros of the $P(z)$-action to $St(GF(q))$. Then, $Saut_{P}$ is a cyclic group of order $q^{k}-1$.
Proof.
Let
$P(z)=a_{0}+a_{1}z+\cdots+a_{k}z^{k}~\mbox{with}~a_{i}\in GF(q)~(i=0,1,\cdots,k),a_{0},a_{k}\neq 0.$
Then, any element in $\Omega_{P}^{q}$ is cyclic with period $q^{k}-1$, since for any $(x_{n}\in GF(q);n\in\mathbb{Z})\in\Omega_{P}^{q}$, $x_{n}$ is determined by $(x_{n-k},x_{n-k+1},\cdots,x_{n-1})$.
Let $\theta_{1},\theta_{2},\cdots,\theta_{k}$ be the roots of $P(z)=0$. Since $P(z)$ is irreducible, they differ each other. Let $\eta$ be a primitive root of $GF(q^{k})=GF(q)(\theta_{1},\cdots,\theta_{k})$. That is, $\eta^{i}=1$ if and only if $i\equiv 0~({\rm mod}~q^{k}-1)$. Then, $\eta$ can be written as $\eta=c_{1}\theta_{1}+\cdots+c_{k}\theta_{1}^{k}$ with $c_{1},\cdots,c_{k}\in GF(q)$. Let
$\lambda_{i}=c_{1}\theta_{i}+\cdots+c_{k}\theta_{i}^{k}~(i=1,\cdots,k).$
Let the matrix $B$ be defined by (16). Then, it is a matrix with entries in $GF(q)$ since the entries are symmetric functions of $\theta_{1},\cdots,\theta_{k}$. Also, it commutes with $C_{P}$, and hence, $B\in Saut_{P}$. Moreover, since $\lambda_{1}=\eta$ has the multiplicative order $q^{k}-1$, we have $B^{i}=I_{k}$ if and only if $i\equiv 0~(mod~q^{k}-1)$. This implies that $Saut_{P}$ is a cyclic group of order $q^{k}-1$ since the cardinality of the set of nonzero $B$ as (16) is at most $q^{k}-1$. ∎
Example 12.
Let us consider $P(z)=z^{2}+1\in GF(3)[z]$. Let $\theta\in GF(9)$ be a root of $z^{2}+1=0$. Then, $\theta+1$ is a primitive roots of $GF(9)$. The following $B$ generates $Saut_{P}$.
$B=\left(\begin{array}[]{cc}2&2\\ \theta&-\theta\end{array}\right)\left(\begin{array}[]{cc}\theta+1&0\\ 0&-\theta+1\end{array}\right)\left(\begin{array}[]{cc}2&2\\ \theta&-\theta\end{array}\right)^{-1}=\left(\begin{array}[]{cc}2&1\\ 2&2\end{array}\right).$
012012012012
Figure 2: $\sigma$-orbits of $\Omega_{P}^{3}$
012012
Figure 3: $B$-orbit of $\Omega_{P}^{3}$
Then $\sigma$-orbits of $\Omega_{P}^{3}$ and the $B$-orbit of $\Omega_{P}^{3}$ are as above.
Acknowledgment: The authors thank Prof. Hajime Kaneko (Tsukuba Univ., Japan) for his useful suggestions. Actually, Theorem 5 is essentially due to him. The first author was supported by TianYuan Visiting Scholar Program of NSFC(Grant No:12426661).